### v7

# A conversion theorem and minimax optimality for continuum contextual bandits

Abstract
We study the contextual continuum bandits problem, where the learner sequentially receives a side information vector and has to choose an action in a convex set, minimizing a function associated to the context. The goal is to minimize all the underlying functions for the received contexts, leading to the contextual notion of regret, which is stronger than the standard static regret. Assuming that the objective functions are $\gamma$-Hölder with respect to the contexts, $0<\gamma\leq 1$, we demonstrate that any algorithm achieving a sub-linear static regret can be extended to achieve a sub-linear contextual regret. We prove a static-to-contextual regret conversion theorem that provides an upper bound for the contextual regret of the output algorithm as a function of the static regret of the input algorithm. We further study implications of this general result for three fundamental cases of dependency of the objective function on the action variable: (a) Lipschitz bandits, (b) convex bandits, (c) strongly convex and smooth bandits. For Lipschitz bandits and $\gamma=1$, combining our results with the lower bound of Slivkins (2014), we prove that the minimax optimal contextual regret for the noise-free adversarial setting scales as $T^{(p+d+1)/(p+d+2)}$ up to logarithmic factors, where $p$ and $d$ are the dimensions of the context and of the action spaces, respectively, and $T$ is the number of queries. Then we prove that, when the evaluations are noisy, the rate of contextual regret in $T$ for convex bandits is the same as for strongly convex and smooth bandits and scales as $T^{(p+\gamma)/(p+2\gamma)}$ up to logarithmic factors.
Lastly, we present a minimax lower bound, implying two key facts. First, obtaining a sub-linear contextual regret may be impossible over functions that are not continuous with respect to the context. Second, for convex bandits and strongly convex and smooth bandits, the algorithms that we propose achieve, up to a logarithmic factor, the minimax optimal rate of contextual regret as a function of the number of queries.
1 Introduction
The continuum bandit problem presents a sequential decision-making challenge where, at each round, an action is taken as a continuous variable followed by suffering a corresponding loss. Various variants of this problem have been explored in the literature. The most extensively studied framework involves the learner receiving a function $f_{t}:\Theta\rightarrow\mathbb{R}$ (random or non-random), at each round $t$, where $\Theta\subseteq\mathbb{R}^{d}$ is a convex body (compact and convex set with a non-empty interior). Based on $\{(y_{k},\boldsymbol{z}_{k})\}_{k=1}^{t-1}$, the learner selects $\boldsymbol{z}_{t}\in\Theta$, suffers loss $f_{t}(\boldsymbol{z}_{t})$, and receives the noisy feedback
$y_{t}=f_{t}(\boldsymbol{z}_{t})+\xi_{t},$
where $\xi_{t}$ denotes the noise variable. The standard goal of the learner is to minimize the static regret
$R_{T}:=\mathbf{E}\Big[\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\min_{\boldsymbol{z}\in\Theta}\sum_{t=1}^{T}f_{t}(\boldsymbol{z})\Big].$ (1)
In the present paper, we assume that the learner has access to side information about the loss function, namely a context at each round. Leveraging this additional information, the learner is expected to minimize another type of regret referred to as the contextual regret. We consider the problem of contextual continuum bandits stated as follows.
Contextual continuum bandits. Let $\Theta\subseteq\mathbb{R}^{d}$ be a convex body. Let $f:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$ be an unknown function. At each round $t$:
  • A context $\boldsymbol{c}_{t}\in[0,1]^{p}$ is revealed by the adversary.
  • Based on the past values $\{(y_{k},\boldsymbol{z}_{k},\boldsymbol{c}_{k})\}_{k=1}^{t-1}$, the learner chooses a query point $\boldsymbol{z}_{t}\in\Theta$ and gets a noisy evaluation of $f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})$:
    $y_{t}=f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})+\xi_{t},$ (2)
    where $\xi_{t}$ is a scalar noise variable.
The learner’s objective in this framework is to minimize the contextual regret, defined as:
$R_{T}^{\mathrm{cntx}}:=\mathbf{E}\Big[\sum_{t=1}^{T}f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\sum_{t=1}^{T}\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\Bigg].$ (3)
Some remarks about this setting are in order. As a general framework for the problem, we assume that the contexts are drawn sequentially from the hypercube $0,1]^{p}$. This presentation is chosen to simplify the notation. However, all the upper bounds on the regret throughout the paper hold if the contexts are drawn from a non-empty compact subset of $0,1]^{p}$. Moreover, at first sight, by introducing the notation $f_{t}(\boldsymbol{z})=f(\boldsymbol{z},\boldsymbol{c}_{t})$, the contextual continuum bandit problem as stated above might appear as a special case of the non-contextual continuum bandit problem with dynamic regret (CBDR). In the CBDR setting, the learner selects the query point $\boldsymbol{z}_{t}$ at time $t$ based on past observations $\{y_{k},\boldsymbol{z}_{k}\}_{k=0}^{t-1}$ to minimize the dynamic regret:
${R_{T}^{\mathrm{dyn}}:=}\mathbf{E}\Big[\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\sum_{t=1}^{T}\min_{\boldsymbol{z}\in\Theta}f_{t}(\boldsymbol{z})\Big].$ (4)
With this notation, we have $R_{T}^{\mathrm{cntx}}=R_{T}^{\mathrm{dyn}}$}. However, the two settings are not equivalent. Indeed, the contextual setting that we consider is “statistically easier" than CBDR since partial information about $f_{t}$ in the form of the context $\boldsymbol{c}_{t}$ is revealed to the learner before the action is chosen. This suggests that a better behavior of the regret is potentially achievable.
Another performance measure for continuum contextual bandits called the $L$-regret was introduced by Hazan and Megiddo (2007):
$R_{T}^{L}:=\mathbf{E}\left[\sum_{t=1}^{T}f_{t}(\boldsymbol{z}_{t})-\min_{\boldsymbol{z}(\cdot)\in X_{L}}\sum_{t=1}^{T}f_{t}(\boldsymbol{z}(\boldsymbol{c}_{t}))\right].$ (5)
Here, the minimum is taken over the class $X_{L}$ of all actions satisfying the Lipschitz condition with respect to the context, with given Lipschitz constant $L$. For $L=0$ the $L$-regret is equal to the static regret $R_{T}$.
Considering $f_{t}(\boldsymbol{z})=f(\boldsymbol{z},\boldsymbol{c}_{t})$ we have that $R_{T}^{\mathrm{cntx}}=R_{T}^{\mathrm{dyn}}\geq R_{T}^{L}\geq R_{T}$. The setting with $L$-regret, where $L>0$, is more challenging than the static regret setting as constant actions trivially satisfy the Lipschitz condition. In turn, controlling the contextual regret $R_{T}^{\mathrm{cntx}}$ is more challenging than the $L$-regret since it evaluates the risk against all possible choices of actions and not only against the Lipschitz actions. By definition, small static regret indicates that the policy is at least as effective as a single constant action performing well on the average of all observed functions during learning. In contrast, small dynamic regret signifies that the policy performs well for each function individually. For non-contextual bandits, the CBDR problem has been recently investigated by Yang et al. (2016); Chen and Giannakis (2018); Zhao et al. (2021); Wei and Luo (2021); Wang (2023) under restrictions on various characteristics of the evolution of $f_{t}$’s over time, which are not interpretable in terms of the smoothness assumptions on $f(\cdot,\cdot)$ used in continuum contextual settings.
In Section 5, we demonstrate that, regardless of the behavior of the function $f(\cdot,\boldsymbol{c})$, if the function $f(\boldsymbol{x},\cdot)$ lacks continuity, obtaining a sub-linear contextual regret may be impossible. Hence, we focus on the setting where $f(\boldsymbol{x},\cdot)$ is a continuous function for any $\boldsymbol{x}\in\mathbb{R}^{d}$. Specifically, in the sequel we adopt the following slightly stronger assumption.
Assumption 1
For $(L,\gamma)\in[0,\infty)\times(0,1]$ we assume that $f\in\mathcal{F_{\gamma}}(L)$, where $\mathcal{F_{\gamma}}(L)$ denotes the class of all functions $f:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$ such that
$|f(\boldsymbol{x},\boldsymbol{c})-f(\boldsymbol{x},\boldsymbol{c}^{\prime})|\leq L\left\lVert\boldsymbol{c}-\boldsymbol{c}^{\prime}\right\rVert^{\gamma},\quad\text{for all}\quad\boldsymbol{x}\in\Theta,\,\,\boldsymbol{c},\boldsymbol{c}^{\prime}\in[0,1]^{p},$ (6)
where $\left\lVert\cdot\right\rVert$ is the Euclidean norm.
In this paper, we develop a meta-algorithm approach to continuum contextual bandit problem. In short it can be described as follows. Let $\mathcal{F}^{\prime}$ be a class of real-valued functions on $\mathbb{R}^{d}$. Assume that we are given an input algorithm $\pi$ achieving a sub-linear static regret in adversarial bandit setting over $T$ runs for the sequence of functions $\{f_{t}(\cdot)\}_{t=1}^{T}$, where each $f_{t}\in\mathcal{F}^{\prime}$. Then, we associate to it an output contextual bandit algorithm with sub-linear contextual regret for any loss $f(\cdot,\cdot)$ belonging to the class $\mathcal{G}_{\gamma}(\mathcal{F}^{\prime}):=\{f\in\mathcal{F_{\gamma}}(L):f(\cdot,\boldsymbol{c})\in\mathcal{F}^{\prime}\ \text{for all}\,\boldsymbol{c}\in[0,1]^{p}\}$. We prove a static-to-contextual regret conversion theorem that provides an upper bound for the contextual regret of the output algorithm as a function of the static regret of the input algorithm. Our construction of the output algorithm consists in discretizing the space of contexts into cells and running an algorithm $\pi$ separately within each cell. Similar constructions can be traced back to Hazan and Megiddo (2007), where the proposed method was not a meta-algorithm valid for any $\pi$ and $\mathcal{F}^{\prime}$ but a single contextual bandit algorithm with a specific input algorithm $\pi$ (projected gradient) and a specific class $\mathcal{F}^{\prime}$ (convex and Lipschitz functions). The analysis given in Hazan and Megiddo (2007) focuses on the $L$-regret and does not provide a control of the contextual regret. The subsequent work Cesa-Bianchi et al. (2017) adopts a similar discretization strategy, with similar limitations. It derives a bound on the $L$-regret when the input algorithm is chosen as Exp3 and the class $\mathcal{F}^{\prime}$ is the set of all 1-Lipschitz functions.
Another approach to static-to-contextual regret conversion is developed in Slivkins (2014). It relies on zooming with dyadic splits and is valid, again, for a specific input class $\mathcal{F}^{\prime}$, namely, a subset of Lipschitz functions, and a specific class of input algorithms $\pi$, namely, those achieving the optimal rate on this particular class. The class of functions $f(\cdot,\cdot)$ considered in Slivkins (2014) is a subclass of $\mathcal{G}_{1}(\mathcal{F}^{\prime})$ satisfying, along with the Lipschitz condition, the additional restrictive condition that for each $\boldsymbol{c}$, there exists $\boldsymbol{z}^{*}(\boldsymbol{c})=\arg\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c})$ such that $\boldsymbol{c}\mapsto f(\boldsymbol{z}^{*}(\boldsymbol{c}),\boldsymbol{c})$ is Lipschitz. The algorithm of Slivkins (2014) requires exact knowledge of the Lipschitz constant, and the resulting bound on the contextual regret is exponential in the dimension of the context.
The contributions of the present paper can be summarized as follows.
  • (i)
    In Section 3, we propose a meta-algorithm that converts any input non-contextual bandit algorithm into an output contextual bandit algorithm. We prove a general conversion theorem that derives the contextual regret of the output algorithm from the static regret of the input algorithm (Theorem 2).
  • (ii)
    We apply our static-to-contextual conversion to handle the following two fundamental settings.
    • (ii-a)
      First, we consider the setting, where $f(\cdot,\cdot)$ is Lipschitz in both arguments, that is, $\gamma=1$ and $\mathcal{F}^{\prime}$ is the class of Lipschitz functions. In this setting, when the noises $\xi_{i}$ are zero, we show that our conversion approach leads to an output algorithm with a rate optimal contextual regret improving upon Slivkins (2014) in several respects. Unlike Slivkins (2014), we do not require the above restrictive additional condition on $f$ nor the knowledge of Lipschitz constant, and our bound scales polynomially in the dimension of the context.
    • (ii-b)
      Second, we consider the setting where $\mathcal{F}^{\prime}$ is the class of all convex functions and the noise in observations (2) is sub-Gaussian. In this case, by applying our conversion theorem together with the static regret bound from Fokkema et al. (2024) we prove that the contextual regret of the proposed output algorithm achieves the minimax optimal rate in $T$ up to a logarithmic factor.
  • (iii)
    In Section 4, we apply our conversion theorem in the setting where $\mathcal{F}^{\prime}$ is the class of strongly convex and smooth functions, and the noise is sub-Gaussian. We first construct a bandit algorithm, see Algorithm 2, which is an extension of the Bandit Convex Optimization (BCO) algorithm from Hazan and Levy (2014) to the problem with noisy queries. We prove a bound on its static regret. Then, combining it with the conversion theorem of Section 3, we derive a bound on the contextual regret of the proposed output algorithm.
  • (iv)
    In Section 5, we establish a minimax lower bound leading to several useful corollaries. First, it implies that the contextual regret bounds from items (ii-b) and (iii) are minimax optimal in $T$ up to logarithmic factors. Second, it shows that if $f(\boldsymbol{x},\cdot)$ lacks continuity, obtaining a sub-linear contextual regret may be impossible.
2 Related work
Contextual bandit problem has been introduced by Woodroofe (1979) in the framework of multi-armed bandits, that is, when the space of actions is finite. The literature on the contextual bandit problem with finite number of arms is now very rich, see, e.g., Goldenshluger and Zeevi (2009); Wang et al. (2005); Bastani and Bayati (2020); Perchet and Rigollet (2013); Rigollet and Zeevi (2010); Pandey et al. (2007); Langford and Zhang (2007); Auer et al. (2002); Kakade et al. (2008); Yang and Zhu (2002); Foster and Rakhlin (2023); Cella et al. (2023, 2022) and the references therein. In this literature, it is not uncommon to assume that the contexts are i.i.d. rather than adversarial.
There was also a considerable work on the problem of contextual bandits with continuous action space: Hazan and Megiddo (2007); Slivkins (2014); Kleinberg et al. (2008); Cesa-Bianchi et al. (2017); Chen and Gallego (2018); Li et al. (2019). These papers deal with adversarial contexts and, with the exception of Li et al. (2019), consider the noiseless setting ($\xi_{i}\equiv 0$). The analysis in Slivkins (2014) addresses the case where the context and the action vary in metric spaces with covering dimensions $p$ and $d$, respectively, under the assumption that $f(\cdot,\cdot)$ is Lipschitz in both arguments and satisfies an additional strong restriction (see the Introduction). The results in Slivkins (2014) provide a lower bound of the order $O(T^{\frac{p+d+1}{p+d+2}})$ on the contextual regret and an algorithm based on zooming techniques that attains this rate up to a logarithmic factor. Bounds for the $L$-regret for continuum contextual bandit algorithms are developed in Hazan and Megiddo (2007); Cesa-Bianchi et al. (2017). The paper Chen and Gallego (2018) deals with the setting where $f(\cdot,\cdot)$ is Lipschitz in both arguments, the actions belong to $[0,1]$, the function $f(\cdot,\boldsymbol{c})$ has a unique minimizer for any context $\boldsymbol{c}$ and exhibits approximately quadratic behavior around the minimizer. Under these assumptions, Chen and Gallego (2018) propose an algorithm with $O(T^{\frac{p+2}{p+4}})$ contextual regret. In Li et al. (2019), the authors addressed the problem of contextual bandits when the functions are strongly convex and smooth and derived a bound for the contextual regret that scales as $O(T^{\frac{p+1}{p+2}})$ up to logarithmic factors. To get this result, Li et al. (2019) additionally assumed that, for each $\boldsymbol{c}\in[0,1]^{p}$, the global minimizer $\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{d}}f(\boldsymbol{x},\boldsymbol{c})$ is an interior point of the constraint set $\Theta$. While this assumption might be hard to check in practice, it does simplify the analysis considerably. In the present paper, we drop this assumption and obtain in Section 4 an upper bound that improves upon the result of Li et al. (2019). Furthermore, in Section 5 we present a matching lower bound demonstrating that the algorithm that we propose is minimax optimal to within a logarithmic factor.
In all the papers on continuum contextual bandits cited above, the contexts are adversarial, so the contextual regret is a variant of the dynamic regret as discussed after (3). The dynamic regret was also in the focus of studies in various settings of online convex optimization. For zero-order optimization (aka gradient-free optimization), the dynamic regret has been studied in Zinkevich (2003); Yang et al. (2016); Zhang et al. (2018); Chen and Giannakis (2018); Zhao et al. (2021). Among these papers, Yang et al. (2016); Chen and Giannakis (2018); Zhao et al. (2021) have addressed the bandit framework, in which the learner is not allowed to query function evaluations outside of the constraint set $\Theta$.
In the framework of adversarial multi-armed contextual bandits with finite number of arms, without imposing any regularity assumptions on the context such as our Assumption 1, Chen et al. (2019); Luo et al. (2018); Auer et al. (2019) achieved sub-linear regret by assuming that the functions $f_{t}$ either change infrequently or vary negligibly over time.
3 A general conversion theorem
In this section, we propose a meta-algorithm that to any input algorithm achieving a sub-linear static regret in adversarial bandit setting associates a contextual bandit algorithm with sub-linear contextual regret. We establish a conversion theorem (Theorem 2 below) that provides an upper bound for the contextual regret of the output algorithm as a function of the static regret of the input algorithm.
Here and in the sequel, unless other assumption is explicitly stated, we assume that the context $\boldsymbol{c}_{t}$ is drawn by an adversary who may be aware of all previous actions of the learner and past information received by the learner. We also assume throughout that for any $\boldsymbol{c}\in[0,1]^{p}$ the minimum $\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c})$ is attained by some $\boldsymbol{z}^{*}(\boldsymbol{c})\in\Theta$.
We consider strategies for choosing the query points such that $\boldsymbol{z}_{0}\in\mathbb{R}^{d}$ is a random variable, and $\boldsymbol{z}_{t}=\pi_{t}(\{\boldsymbol{c}_{k},\boldsymbol{z}_{k},y_{k}\}_{k=1}^{t-1},\boldsymbol{c}_{t},\boldsymbol{\zeta}_{t})$ for $t\geq 1$, where $\pi_{t}$ is a measurable function with values in $\mathbb{R}^{d}$, and $(\boldsymbol{\zeta}_{t})_{t\geq 1}$ is a sequence of random variables with values in a measurable space $(\mathcal{Z},\mathcal{U})$ such that $\boldsymbol{\zeta}_{t}$ is independent of $\{\boldsymbol{c}_{k},\boldsymbol{z}_{k},y_{k}\}_{k=1}^{t-1},\boldsymbol{c}_{t}$. We call the process $\pi=(\pi_{t})_{t\geq 1}$ a randomized procedure and we denote the set of all randomized procedures by $\Pi$.
Let $\mathcal{F}^{\prime}$ be a class of real-valued functions on $\mathbb{R}^{d}$. We define $\mathcal{F}=\{f:f(\cdot,\boldsymbol{c})\in\mathcal{F}^{\prime}\text{ for all }\boldsymbol{c}\in[0,1]^{p}\}$. Assume that there exists a randomized procedure $\pi$, for which we can control its static regret over $T$ runs for the sequence of functions $\{f(\cdot,\boldsymbol{c}_{t})\}_{t=1}^{T}$ where $f\in\mathcal{F}$. Namely, if $\{\boldsymbol{z}_{t}\}_{t=1}^{T}$ are outputs of $\pi$, then there exist $F:\mathbb{R}^{+}\cup\{0\}\to\mathbb{R}^{+}\cup\{0\}$, $F_{1}:\mathbb{R}^{+}\cup\{0\}\to\mathbb{R}^{+}\cup\{0\}$ such that $F$ is a concave function, $F_{1}$ is a non-decreasing function, and
${\sup_{f\in\mathcal{F}}\,}\mathbf{E}\left[\sum_{t=1}^{T}f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}\sum_{t=1}^{T}f(\boldsymbol{z},\boldsymbol{c}_{t})\right]\leq F(T){F_{1}(T)}.$ (7)
In particular, we will be interested in the case where there exists $(\tau_{1},\tau_{2},\tau_{3},\mathrm{T}_{0})\in[0,\infty)\times[0,1)\times[0,\infty)\times[1,\infty)$ such that, for $T\geq\mathrm{T}_{0}$,
$F(T)\leq Cd^{\tau_{1}}T^{\tau_{2}}\quad\text{and}\quad F_{1}(T)\leq\log^{\tau_{3}}(T+1),$ (8)
where $C>0$ is a constant that does not depend on $T$ and $d$.
Definition 1
We call a randomized procedure that satisfies (7) and (8) a $(\mathcal{F},\tau_{1},\tau_{2},\tau_{3},\mathrm{T}_{0})$-consistent randomized procedure and we denote the set of all such procedures by $\Pi(\mathcal{F},\tau_{1},\tau_{2},\tau_{3},\mathrm{T}_{0})$.
Input: Randomized procedure $\pi=(\pi_{t})_{t=1}^{\infty}$, parameter $K\in\mathbb{N}$ and a partition $\{B_{i}\}_{i=1}^{K^{d}}$ of $0,1]^{p}$
Initialization: $N_{1}(0),\dots,N_{K^{p}}(0)=0$, $H(1),\dots,H(K^{p})=\{0\},\text{ and }\boldsymbol{z}_{0}=\boldsymbol{0}$
for $t=1,\ldots,T$ do
 if $\boldsymbol{c}_{t}\in B_{i}$ then
     2mm Let $N_{i}(t)=N_{i}(t-1)+1$
    // Increment
     2mm Based on $\boldsymbol{c}_{t}$ and $\{y_{k},\boldsymbol{z}_{k},\boldsymbol{c}_{k}\}_{k\in H(i)}$, use $\pi_{N_{i}(t)}$ to choose $\boldsymbol{z}_{t}\in\Theta$
    // Choosing the query point
     2mm $y_{t}=f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})+\xi_{t}$
    // Query
    $H(i)=H(i)\cup\{t\}$
    // Update the set of indices
   end if
end for
Algorithm 1
Algorithm 1 is a meta-algorithm that assures regret conversion. It uses a partition $\{B_{i}\}_{i=1}^{K^{p}}$ of the hypercube $0,1]^{p}$, where all the cells $B_{i}$ are hypercubes of equal volume with edges of length $1/K$. Fix a consistent randomized procedure $\pi$ for the class of functions $\mathcal{F}$. At each round $t$, if the context $\boldsymbol{c}_{t}$ falls into the cell $B_{i}$, Algorithm 1 chooses the query point $\boldsymbol{z}_{t}$ by the procedure $\pi$ based only on the past contexts from the same cell $B_{i}$ and the corresponding past query points. Since no more than $T$ cells $B_{i}$ can be visited and need to be memorized Algorithm 1 is polynomial time provided the input algorithm $\pi$ is polynomial time.
Let $\{\boldsymbol{c}_{i_{j}}\}_{j=1}^{N_{i}(T)}$ be all the contexts belonging to $B_{i}$ revealed up to round $T$, where $N_{i}(T)$ is the total number of such contexts, and let $\{\boldsymbol{z}_{i_{j}}\}_{j=1}^{N_{i}(T)}$ be the outputs of $\pi$ restricted to the contexts that belong to $B_{i}$ (the inner loop of Algorithm 1). This segment of the algorithm involves saving the index of the rounds associated with the received contexts in cell $B_{i}$. The set $H(i)$ is the collection of such indices. Denote by $\boldsymbol{b}_{i}$ the barycenter of $B_{i}$, and let $\boldsymbol{x}_{i}^{*}\in\operatorname*{arg\,min}_{\boldsymbol{x}\in\Theta}f(\boldsymbol{x},\boldsymbol{b}_{i})$. Then,
$\mathbf{E}\bigg[\sum_{t=1}^{T}\bigg(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},$ $\boldsymbol{c}_{t})\bigg)\bigg]=\mathbf{E}\bigg[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}_{i}^{*},\boldsymbol{c}_{i_{j}})+\Delta_{i_{j}}\right)\bigg]$
$\leq\mathbf{E}\bigg[\sum_{i=1}^{K^{p}}\big(\sum_{j=1}^{N_{i}(T)}f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-\min_{\boldsymbol{z}\in\Theta}\sum_{j=1}^{N_{i}(T)}f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})+\sum_{j=1}^{N_{i}(T)}\Delta_{i_{j}}\big)\bigg],$
where $\Delta_{i_{j}}=f(\boldsymbol{x}_{i}^{*},\boldsymbol{c}_{i_{j}})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})$. One may notice that $\Delta_{i_{j}}\leq 2\max_{\boldsymbol{z}\in\Theta}|f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{z},\boldsymbol{b}_{i})|$ (see the proof of Theorem 2), so that under Assumption 1 we have $\Delta_{i_{j}}\leq 2L(\sqrt{p}/K)^{\gamma}$. Thus, we can control the dynamic regret by the sum of the static regret, which grows with $K$, and the bias term comprised of $\Delta_{i_{j}}$’s, which decreases with $K$. There exists an optimal $K$ exhibiting a trade-off between the static regret and the bias. In Theorem 2, we show that if $K$ is tuned optimally,
$K=K(\tau_{1},\tau_{2},\tau_{3},\gamma)\triangleq\max\left(1,\left\lfloor\left(Lp^{\frac{\gamma}{2}}d^{-\tau_{1}}{T^{(1-\tau_{2})}\log^{-\tau_{3}}(T+1)}\right)^{\frac{1}{p(1-\tau_{2})+\gamma}}\right\rfloor\right),$ (9)
then Algorithm 1 achieves a sub-linear dynamic regret.
Theorem 2
Let $(L,\gamma)\in[0,\infty)\times(0,1]$. Let $\pi$ be a randomized procedure such that (7) holds with a concave function $F$, a non-decreasing function $F_{1}$, and let $\boldsymbol{z}_{t}$’s be the updates of Algorithm 1 with a positive integer $K$. Then,
$\sup_{f\in\mathcal{F}\cap\mathcal{F}_{\gamma}(L)}\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\big)\big]\leq K^{p}F\left(\frac{T}{K^{p}}\right){F_{1}(T)}+2LT\left(\frac{\sqrt{p}}{K}\right)^{\gamma}.$ (10)
If, in addition, $\pi\in\Pi(\mathcal{F},\tau_{1},\tau_{2},\tau_{3},\mathrm{T}_{0})$ for a triplet $(\tau_{1},\tau_{2},\mathrm{T}_{0})\in[0,\infty)\times[0,1)\times[1,\infty)$, and $K=K(\tau_{1},\tau_{2},\tau_{3},\gamma)$ (see (9)), then for any $T\geq K^{p}\mathrm{T}_{0}$, we have
$\sup_{f\in\mathcal{F}\cap\mathcal{F}_{\gamma}(L)}\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\big)\big]$
$\hskip 28.45274pt\leq C^{\prime}\max\bigg(d^{\tau_{1}}T^{\tau_{2}}L_{1},\left(Lp^{\frac{\gamma}{2}}\right)^{\frac{p(1-\tau_{2})}{p(1-\tau_{2})+\gamma}}\left(d^{\tau_{1}}L_{1}\right)^{\frac{\gamma}{p(1-\tau_{2})+\gamma}}T^{\frac{(p-\gamma)(1-\tau_{2})+\gamma}{p(1-\tau_{2})+\gamma}}\bigg).$ (11)
where $L_{1}=\log^{\tau_{3}}(T+1)$ and $C^{\prime}>0$ is a constant independent of $T$, $d$, and $p$.
Note that the Lipschitz constant $L$ measures how sensitive the function $f$ is with respect to variations of the contexts. If $L=0$ the static and contextual regrets are equivalent, and it is not surprising that (8) and (2) have the same rates. We also note that a bound that depends in the same way as (2) on the parameters $p$, $d$, and $T$ can be achieved with a choice of $K$ independent of $L$. In deriving (2) we assumed that the learner knows the value of $L$ but this is only done to illustrate the above remark.
The approach to contextual bandits based on discretizing the space of contexts and running some algorithms $\pi$ separately for each discretization item is quite common, see Hazan and Megiddo (2007); Kleinberg et al. (2008); Bubeck et al. (2008); Slivkins (2014); Cesa-Bianchi et al. (2017); Li et al. (2019). With the exception of Slivkins (2014), each of these papers, as well as follow-up works based on similar arguments, explored a specific algorithm $\pi$, a specific class $\mathcal{F}$, and provided bounds on the contextual regret (or on $L$-regret) using their particular properties. More generally, (Slivkins, 2014, Theorem 21) provides a static-to-contextual regret conversion device, where the class $\mathcal{F}$ is still quite specific (defined by conditions (S1) and (S2) below) but the input algorithm $\pi$ is not fixed and allowed to belong to the class of algorithms with $\tilde{O}(T^{\frac{1}{p+2}})$ static regret. The bound on the contextual regret in (Slivkins, 2014, Theorem 21) includes the doubling dimension, which is exponential in the dimension of the context. In contrast, Theorem 2 is a conversion result valid for a general class $\mathcal{F}$, a general input algorithm $\pi$, and its bound is free from exponential factors. It yields a ready-to-use tool for exploring various particular settings. Below we demonstrate its application to the three major contextual settings, where the class $\mathcal{F}$ is defined by Assumption 1 combined with one of the following assumptions on the class of functions $f(\cdot,\boldsymbol{c})$ for any fixed $\boldsymbol{c}$:
  • (a)
    $\mathcal{F}^{\prime}$ is a class of Lipschitz functions,
  • (b)
    $\mathcal{F}^{\prime}$ is the class of all convex functions,
  • (c)
    $\mathcal{F}^{\prime}$ is a class of strongly convex and smooth functions.
For the sake of comparison, we recall that the class of loss functions $\mathcal{F}$ considered in (Slivkins, 2014, Theorem 21) is defined by the following two conditions:
00 (S1) $f(\boldsymbol{z},\boldsymbol{c})$ is Lipschitz in $(\boldsymbol{z},\boldsymbol{c})$ with Lipschitz constant 1,
00 (S2) for each context $\boldsymbol{c}$, there exists $\boldsymbol{z}^{*}(\boldsymbol{c})\in\operatorname*{arg\,min}_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c})$ such that the map $\boldsymbol{c}\mapsto f(\boldsymbol{z}^{*}(\boldsymbol{c}),\boldsymbol{c})$ is Lipschitz, with Lipschitz constant 1.
0 Condition (S2) does not follow from (S1) and requires some strong additional properties of $f$ that are not detailed in Slivkins (2014). Moreover, as emphasized in Slivkins (2014) the assumption that the Lipschitz constant is 1 can be relaxed but in that case the algorithm should know the Lipschitz constant.
For the rest of this section, we focus on the implications of Theorem 2 for two families of functions corresponding to items (a) and (b) above. Observe that considering (a) contains as a special case the class $\mathcal{F}$ of functions $f(\boldsymbol{z},\boldsymbol{c})$ that are Lipschitz in $(\boldsymbol{z},\boldsymbol{c})$ with no further restriction. We do not need the hardly interpretable additional condition (S2), nor the knowledge of the Lipschitz constant.
(a) Lipschitz bandits. Lipschitz non-contextual bandits have been studied extensively in the literature, cf. Lattimore and Szepesvári (2020) and the references therein. The main idea is to reduce the Lipschitz bandit problem to a multi-armed bandit problem with finite number of arms, often employing zooming techniques with dyadic splits; see, for example, Kleinberg (2004); Kleinberg et al. (2008); Bubeck et al. (2008). While these methods benefit from near optimal in $T$ static regret (of the order $\tilde{O}(T^{\frac{d+1}{d+2}})$), the regret bounds depend exponentially on the dimension $d$. A recent improvement due to Podimata and Slivkins (2021) shows that polynomial dependence on the dimension is achievable. In Podimata and Slivkins (2021), the authors proposed an algorithm with static regret that scales as $d^{\frac{1}{2}}T^{\frac{d+1}{d+2}}\log^{5}(T)$ for $\Theta=[0,1]^{d}$ (cf. (Podimata and Slivkins, 2021, Theorem 2) with $z=d$ for the worst-case scenario). Taking their method as an input $\pi$ of our meta-algorithm and combining the bound of (Podimata and Slivkins, 2021, Theorem 2) with Theorem 2 we obtain the following corollary.
Corollary 3
Assume that $\Theta=[0,1]^{d}$ and that the observations in (2) are noise-free, that is, $\xi_{t}=0$ for all positive integers $t$. Let $\mathcal{F}$ be the set of functions such that $|f(\boldsymbol{x},\boldsymbol{c})-f(\boldsymbol{x}^{\prime},\boldsymbol{c})|\leq L^{\prime}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in[0,1]^{d}$, $\boldsymbol{c}\in[0,1]^{p}$, where $L^{\prime}>0$ is a constant. Let $\boldsymbol{z}_{t}$’s be the updates of Algorithm 1 with the input algorithm $\pi$ taken as the Lipschitz bandit algorithm from Podimata and Slivkins (2021). Then, under the optimal choice of $K$ we have
$\sup_{f\in\mathcal{F}\cap\mathcal{F}_{\gamma}(L)}\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})$ $\big)\big]\leq C\max\bigg(d^{1/2}T^{(d+1)/(d+2)}L_{1},$
$\big(p^{\gamma p/2}\big(d^{\frac{1}{2}}L_{1}\big)^{\gamma(d+2)}T^{p+(d+1)\gamma}\big)^{1/(p+(d+2)\gamma)}\bigg),$
where $L_{1}=\log^{5}(T+1)$, and $C>0$ is a constant that does not depend on $T$, $d$, and $p$.
Proof By (Podimata and Slivkins, 2021, Theorem 2) the input algorithm $\pi$ satisfies the static regret bound (7) with
$F(T)\leq C_{1}d^{1/2}T^{(d+1)/(d+2)},\quad\text{and}\quad F_{1}(T+1)\leq\log^{5}(T+1),$
where $C_{1}>0$ is a constant independent of $T$, $d$, and $p$. Thus, we can apply Theorem 2 with $\tau_{1}=1/2$, $\tau_{2}=(d+1)/(d+2)$, and $\tau_{3}=5$, which yields the result.
In particular, in the case $\gamma=1$, that is, when $f$ is Lipschitz continuous in both arguments, Corollary 3 establishes the rate in $T$ as $T^{(p+d+1)/(p+d+2)}$ up to a logarithmic factor and a weak factor depending on $p,d$, which does not exceed $O(\sqrt{pd})$. On the other hand, for $\gamma=1$, Slivkins (2014) provides a matching lower bound on the contextual regret of the same order $T^{(p+d+1)/(p+d+2)}$. Together with Corollary 3, it implies that the rate $T^{(p+d+1)/(p+d+2)}$ is minimax optimal as function of $T$, up to logarithmic factors, and our algorithm attains the minimax rate. Notice that the minimax optimality of this rate on the class of functions $f$ that are Lipschitz continuous in both arguments was not established in Slivkins (2014), contrary to what is sometimes claimed. Indeed, Slivkins (2014) provided an upper bound on a substantially smaller class where additionally condition (S2) holds. Furthermore, the upper bound in (Slivkins, 2014, Theorem 21) is proportional to doubling dimension, which scales exponentially in $p$, while Corollary 3 involves only a weak dimension dependent factor that never exceeds $O(\sqrt{pd})$.
(b) Convex bandits. The study of non-contextual bandit convex optimization was initiated by Flaxman et al. (2004) and Kleinberg (2004). There exists an extensive literature on this problem for both stochastic and adversarial settings, see Orabona (2019); Lattimore and Szepesvári (2020); Lattimore (2024) and the references therein. Bounds on the static regret in adversarial setting were developed in Bubeck et al. (2017); Bubeck and Eldan (2018); Lattimore and Szepesvári (2020); Fokkema et al. (2024). Papers Bubeck et al. (2017); Bubeck and Eldan (2018); Lattimore and Szepesvári (2020) considered the setting with no noise ($\xi_{t}=0$ for all $t$) providing successive improvements of static regret bounds. The best known result Lattimore and Szepesvári (2020) yields a bound of the order $d^{2.5}\sqrt{T}$ derived as an evaluation of the minimax regret without proposing an explicit algorithm (here and below we ignore the logarithmic factors). An important progress was recently achieved in Fokkema et al. (2024), where the authors proposed a polynomial time algorithm attaining a regret of the order $d^{3.5}\sqrt{T}$ in adversarial setting with sub-Gaussian noise, cf. (Fokkema et al., 2024, Theorem 1). Combining this result, which is the state of the art, with Theorem 2 we obtain Corollary 4 below for the contextual convex bandit setting. As (Fokkema et al., 2024, Theorem 1) does not explicitly detail the dependence on logarithmic factors, we adopt the same convention and present the bound up to logarithmic factors.
Corollary 4
Let $\mathcal{F}$ be the set of functions $f$ such that $f(\cdot,\boldsymbol{c})$ is convex for all $\boldsymbol{c}\in[0,1]^{p}$. Assume that, conditionally on the past observations, $\xi_{t}$ in (2) is a zero-mean $\sigma$-sub-Gaussian random variable, that is, for any $\lambda\in{\mathbb{R}}$ and some $\sigma>0$ we have $\mathbf{E}\left[\exp\left(\lambda\xi_{t}\right)|\{\boldsymbol{c}_{k},\boldsymbol{z}_{k},y_{k}\}_{k=1}^{t-1},\boldsymbol{c}_{t}\right]\leq\exp\left(\sigma^{2}\lambda^{2}/2\right)$ [^1] Requiring the inequality for all $\lambda\in{\mathbb{R}}$ grants that $\xi_{t}$ has zero conditional mean. for all positive integers $t$. Let $\boldsymbol{z}_{t}$’s be the updates of Algorithm 1 with the input algorithm $\pi$ taken as the convex bandit algorithm from Fokkema et al. (2024). Then, under the optimal choice of $K$ we have
$\sup_{f\in\mathcal{F}\cap\mathcal{F}_{\gamma}(L)}\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})$ $\big)\big]\leq C\max\left(d^{7/2}\sqrt{T},p^{\gamma p/(2p+4\gamma)}d^{7\gamma/\left(p+2\gamma\right)}T^{\frac{p+\gamma}{p+2\gamma}}\right),$
where $C>0$ is a factor that depends polynomially on $\log(T)$ and $\log(d)$ and does not depend on $p$.
Proof By (Fokkema et al., 2024, Theorem 1) the input algorithm $\pi$ satisfies the static regret bound (7) with $F(T)\leq C^{\prime}d^{7/2}\sqrt{T}$ and $F_{1}(T)=\log^{a}(T)$ for some $a>0$ and a constant $C^{\prime}>0$ that depends polynomially on $\log(d)$ and does not depend on $p$. Thus, we can apply Theorem 2 with $\tau_{1}=7/2$ and $\tau_{2}=1/2$. Taking $K$ as defined in (9) with $\tau_{1}=7/2$, $\tau_{2}=1/2$, and $\tau_{3}=a$ we obtain the result of the theorem (with the dependency with respect to logarithmic factors hidden in $C$).
Thus, for convex contextual bandits we have an algorithm with the contextual regret that scales in $T$ as $T^{\frac{p+\gamma}{p+2\gamma}}$ up to a logarithmic factor. The dependence on $d$ and $p$ is mild. It is given by a factor that never exceeds $p^{1/2}d^{7/3}$, to within a power of $\log(d)$. Combining this result with the lower bound of Theorem 11 (see Section 5 below) we conclude that the minimax optimal rate in $T$ for convex contextual bandits is of the order $T^{\frac{p+\gamma}{p+2\gamma}}$ up to a logarithmic factor.
4 Strongly convex and smooth functions
In this section, we apply the regret conversion approach developed in Section 3 to the contextual bandits problem with the objective functions $f$ such that $f(\cdot,\boldsymbol{c})$ is $\beta$-smooth and $\alpha$-strongly convex for any $\boldsymbol{c}\in[0,1]^{p}$.
Definition 5
Let $\beta>0$. Function $g:\mathbb{R}^{d}\to\mathbb{R}$ is called $\beta$-smooth, if it is continuously differentiable and for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$ it satisfies
$\left\lVert\nabla g(\boldsymbol{x})-\nabla g(\boldsymbol{y})\right\rVert\leq\beta\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert.$
Definition 6
Let $\alpha>0$. A differentiable function $g:\mathbb{R}^{d}\to\mathbb{R}$ is called $\alpha$-strongly convex, if for any $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{d}$ it satisfies
$g(\boldsymbol{x})\geq g(\boldsymbol{y})+\langle\nabla g(\boldsymbol{y}),\boldsymbol{x}-\boldsymbol{y}\rangle+\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert^{2}.$
We impose the following assumption on the function $f$.
Assumption 2
For $(\alpha,\beta,M)\in(0,\infty)\times[\alpha,\infty)\times(0,\infty)$, we assume that $f\in\mathcal{F}_{\alpha,\beta}(M)$, where $\mathcal{F}_{\alpha,\beta}(M)$ denotes the class of all functions $f:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$ such that for all $\boldsymbol{c}\in[0,1]^{p}$ the following holds:
  • i)
    $f(\cdot,\boldsymbol{c})$ is $\alpha$-strongly convex.
  • ii)
    $f(\cdot,\boldsymbol{c})$ is $\beta$-smooth.
  • iii)
    $\max_{\boldsymbol{x}\in\Theta}|f(\boldsymbol{x},\boldsymbol{c})|\leq M$.
The non-contextual bandit setting with strongly convex and smooth functions has been investigated in Hazan and Levy (2014) and Ito (2020). Both Hazan and Levy (2014) and Ito (2020) studied the problem with no noise ($\xi_{t}\equiv 0$ for all $t$). In Hazan and Levy (2014), the authors proposed an algorithm leveraging an interior point method and a self-concordant function. They achieved a static regret, which is nearly optimal, ranging from $d\sqrt{T}$ to $d^{3/2}\sqrt{T}$ depending on the geometry of the constraint set. Assuming that all minimizers of the objective functions are in the interior of the constraint set, Ito (2020) introduced a procedure that achieves static regret of the optimal order $d\sqrt{T}$ (however, with no polynomial time guarantee). In this section, we extend the algorithm of Hazan and Levy (2014) to the setting with noisy function evaluation as in (2) and derive a bound on the static regret of the proposed algorithm. Then, using it as an input algorithm $\pi$ of our meta-algorithm we obtain a bound on the contextual regret of the resulting procedure.
Following the approach of Section 3, we first construct a consistent randomized procedure, that is, a procedure with sub-linear static regret, cf. (7). To this end, our starting point is the BCO algorithm introduced in Hazan and Levy (2014). The BCO algorithm makes use of the interior point method, which is a vital tool in convex optimization, see, e.g, Nesterov and Nemirovskii (1994). A pioneering role of introducing this method to the field of online learning is credited to Abernethy et al. (2008). We cannot directly apply the BCO algorithm as defined in Hazan and Levy (2014) since Hazan and Levy (2014) deals with the setting, where the observations $y_{t}$ are noise-free. Therefore, we propose a modification of the BCO algorithm adapted to noisy observations (2) with sub-Gaussian noise $\xi_{t}$. It is defined in Algorithm 2. In what follows, $[T]$ denotes the set of all positive integers less than or equal to $T$. We denote the unit Euclidean ball in $\mathbb{R}^{d}$ by $B^{d}$ and the unit Euclidean sphere by $\partial B^{d}$.
Assumption 3
There is $\sigma>0$ such that for all $t\in[T]$ the following holds: (i) $\xi_{t}$ is $\sigma$-sub-Gaussian, i.e., for any $\lambda>0$ we have $\mathbf{E}\left[\exp\left(\lambda\xi\right)\right]\leq\exp\left(\sigma^{2}\lambda^{2}/2\right)$; (ii) $\xi_{t}$ is independent of $\boldsymbol{\zeta}_{t}$.
We emphasize that the mutual independence of the noise variables $\xi_{t}$ is not assumed.
The main ingredient of the interior point method is a self-concordant barrier defined as follows.
Definition 7
Let $\text{int}(\Theta)$ be the interior of the set $\Theta$. For $\mu>0$, a function $\mathcal{R}:\text{int}(\Theta)\to\mathbb{R}$ is a self-concordant function if the following holds.
  • $\mathcal{R}$ is three times differentiable and converges to infinity for any trajectory that converges to the boundary of $\Theta$, i.e., $\mathcal{R}(\boldsymbol{y})\to\infty$ as $\boldsymbol{y}\to\partial\Theta$.
  • For any $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\boldsymbol{y}\in\text{int}(\Theta)$ we have
    $|\nabla^{3}\mathcal{R}(\boldsymbol{y})[\boldsymbol{x},\boldsymbol{x},\boldsymbol{x}]|\leq 2\left(\nabla^{2}\mathcal{R}(\boldsymbol{y})[\boldsymbol{x},\boldsymbol{x}]\right)^{\frac{3}{2}},$
    where for $m\in{1,2,3}$, and $\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}\in\mathbb{R}^{d}$
    $\nabla^{m}\mathcal{R}(\boldsymbol{y})[\boldsymbol{v}_{1},\dots,\boldsymbol{v}_{m}]=\frac{\partial^{m}}{\partial t_{1}\dots\partial t_{m}}\mathcal{R}\left(\boldsymbol{y}+t_{1}\boldsymbol{v}_{1}+\dots+t_{m}\boldsymbol{v}_{m}\right)|_{t_{1}=0,\dots,t_{m}=0}.$
If, in addition, for some $\mu>0$, and any $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\boldsymbol{y}\in\text{int}(\Theta)$ we have
$|\nabla\mathcal{R}(\boldsymbol{y})[\boldsymbol{x}]|\leq\mu^{\frac{1}{2}}\left(\nabla^{2}\mathcal{R}(\boldsymbol{y})[\boldsymbol{x},\boldsymbol{x}]\right)^{\frac{1}{2}},$
then the function $\mathcal{R}$ is called a $\mu$-self-concordant barrier.
Since $\Theta$ is bounded, $\nabla^{2}\mathcal{R}(\boldsymbol{x})$ is positive definite, the minimizer of $\mathcal{R}$ in $\text{int}(\Theta)$ is unique, see (Nemirovski, 2004, Chapter 3, Section V), and $\boldsymbol{x}_{t}$ in Algorithm 2 is uniquely defined as well. Moreover, without loss of generality, we assume that $\min_{\boldsymbol{x}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{x})=0$. Self-concordant barriers can be chosen depending on the geometry of $\Theta$. For the examples of $\mathcal{R}$ and the corresponding values of $\mu$, we refer the reader to Appendix D.
Input: $\mu$-self-concordant barrier function $\mathcal{R}$ on $\Theta$, and $q_{T}=M+2\sigma\sqrt{\log(T+1)}$
Initialization: $\boldsymbol{x}_{0}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{x})$
for $t=1,\ldots,T$ do
 $\eta_{t}=(4dq_{T})^{-1}\min\left(1,\sqrt{\frac{16(\mu+\beta/\alpha)\log(t+1)}{t}}\right)$
 // Update the step size
 $P_{t}=\left(\nabla^{2}\mathcal{R}(\boldsymbol{x}_{t-1})+\eta_{t}\alpha t\mathbf{I}_{d}\right)^{-\frac{1}{2}}$
 // Update the perturbation parameter
   Generate $\boldsymbol{\zeta}_{t}$ uniformly distributed on $\partial B^{d}$
 // Randomization
 $\boldsymbol{z}_{t}=\boldsymbol{x}_{t-1}+P_{t}\boldsymbol{\zeta}_{t}$
 // Output
 $y_{t}=f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})+\xi_{t}\quad$
 // Query
 $\boldsymbol{g}_{t}=dy_{t}P_{t}^{-1}\boldsymbol{\zeta}_{t}$
 // Estimate the gradient
 $\boldsymbol{x}_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\Theta}\eta_{t}\sum_{k=1}^{t}\left(\langle\boldsymbol{g}_{k},\boldsymbol{x}\rangle+\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{k-1}\right\rVert^{2}\right)+\mathcal{R}(\boldsymbol{x})$
 // Iterate
end for
Algorithm 2
The difference between Algorithm 2 and the BCO algorithm of Hazan and Levy (2014) is in the assigned step size. In Hazan and Levy (2014) the step size is constant and depends on the horizon $T$ while it varies with $t$ in Algorithm 2. Next, in Algorithm 2, the step size is downscaled by $q_{T}=M+2\sigma\sqrt{\log(T+1)}$ whereas in Hazan and Levy (2014) it is downscaled by $M$ without any additional logarithmic term. This difference is due to the fact that the observations in (2) are noisy whereas in Hazan and Levy (2014) they are assumed to be noise-free. In our analysis, we rely on the fact that $\eta_{t}\left\lVert P_{t}\boldsymbol{g}_{t}\right\rVert$ is uniformly bounded with high probability. To ensure this, it is enough to scale down the step size by adding the above logarithmic term to $M$. Indeed, due to Assumption 3, we have that $\max_{t\in[T]}|y_{t}|\leq q_{T}$ with a suitably high probability (see the details in the proof of Theorem 8).
Algorithm 2 algorithm selects $\boldsymbol{z}_{t}\in\{\boldsymbol{x}_{t-1}+P_{t}\boldsymbol{\zeta}|\boldsymbol{\zeta}\in\partial B^{d}\}\subseteq D\triangleq\{\boldsymbol{z}|(\boldsymbol{z}-\boldsymbol{x}_{t-1})^{\top}\nabla^{2}\mathcal{R}(\boldsymbol{x}_{t-1})(\boldsymbol{z}-\boldsymbol{x}_{t-1})\leq 1\}$, in which $D$ is the Dikin ellipsoid. Consequently, $D\subseteq\Theta$; see (Nemirovski, 2004, Proposition 14.2.2) for a proof. Therefore, $\boldsymbol{z}_{t}\in\Theta$ for all $t\in[T]$.
Theorem 8
Let Assumptions 2 and 3 hold. Then, the static regret of Algorithm 2 satisfies
$\mathbf{E}\left[\sum_{t=1}^{T}f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}\sum_{t=1}^{T}f(\boldsymbol{z},\boldsymbol{c}_{t})\right]\leq F(T)F_{1}(T),$
where $F(\cdot)$ is a concave function defined for $x\geq 0$ by
$F(x)=11\sqrt{\nu x\log(x+1)}+\nu\left(\log(x+1)+2\log(\nu+1)\right),$
and $F_{1}(\cdot)$ is a non-decreasing function such that for $x\geq 0$
$F_{1}(x)=\frac{d}{4}\left(5.3M+2\sigma\sqrt{\log(x+1)}\right),$
with $\nu=16(\mu+\beta/\alpha)$. Moreover, for $T\geq\nu^{2}/16$, we have
$F(T)\leq 17\sqrt{\nu T\log(T+1)},$ (12)
and subsequently
$F(T)F_{1}(T)\leq 17d\left(5.3M+2\sigma\sqrt{\log(T+1)}\right)\left(\left(\mu+\frac{\beta}{\alpha}\right)T\log(T+1)\right).$ (13)
For the case of noise-free observations in (2) we have $\sigma=0$, and (12) has the same order as the bound on the regret provided in Hazan and Levy (2014). We also note that, for the case of bounded noise ($|\xi_{t}|\leq\sigma$), $q_{T}$ can be set as $q_{T}=M+\sigma$ and a refined regret in (12) can be derived, where the term $2\sigma\sqrt{\log(T+1)}$ is replaced by $\sigma$. Finally, even for the case of no noise $\sigma=0$ we improve on the result of Hazan and Levy (2014) since, unlike Hazan and Levy (2014), our Algorithm 2 is anytime and does not require the knowledge of $T$ when $\sigma=0$.
It is a known fact (see Appendix D) that, for any convex body $\Theta$, one can construct a $\mu$-self-concordant barrier such that for $\rho\in[0,1]$, $\mu\leq\texttt{A}d^{\rho}$, where $\texttt{A}>0$ is a constant depending only on $\Theta$ and independent of $d$. This fact and Theorem 8 imply that $\Pi\left(\mathcal{F}_{\alpha,\beta}(M),1+\rho/2,1/2,1,16(\mu+\beta/\alpha)^{2}\right)$ is non-empty. In Shamir (2013); Akhavan et al. (2020, 2024), the authors established a lower bound of the order of $d/\sqrt{T}$ for the optimization error when the noise is Gaussian (Shamir (2013)) or satisfies more general assumptions (Akhavan et al. (2020, 2024)). By convexity, this lower bound, when multiplied by $T$, provides a lower bound for the static regret in the case $L=0$, that is, when $f(\boldsymbol{x},\boldsymbol{c}_{1})=f(\boldsymbol{x},\boldsymbol{c}_{t})$ for all $\boldsymbol{x}\in\mathbb{R}^{d}$ and $t\in[T]$. Thus, the lower bound in Shamir (2013); Akhavan et al. (2020, 2024) results in a valid lower bound of the order of $d\sqrt{T}$ for the static regret in the noisy setting. Therefore, if $\rho=0$, which is the case when $\Theta$ is a Euclidean ball, (12) is optimal up to a logarithmic factor. In the worst case scenario ($\rho=1$), for instance, when $\Theta$ is the $\ell_{\infty}$-ball or the simplex (see Appendix D), akin to Hazan and Levy (2014), our algorithm exhibits an optimality gap of at most $d^{1/2}$ up to logarithmic factors.
In order to enable the regret conversion via Theorem 2, we need Assumption 1, that is, the premise that $f$ belongs to $\mathcal{F}_{\gamma}(L)$. Since it always holds that $\mu\leq\texttt{A}d^{\rho}$, where $\rho\in[0,1]$ and $\texttt{A}>0$ is a numerical constant, we obtain the following corollary of Theorems 2 and 8.
Input: Parameter $K\in\mathbb{N}$ and a partition of $0,1]^{p}$ namely $(B_{i})_{i=1}^{K^{p}}$, $\mu$-self-concordant barrier $\mathcal{R}$
Initialization: $N_{1}(0),\dots,N_{K^{p}}(0)=0$, $H(1),\dots,H(K^{p})=\{0\}$, $\boldsymbol{x}_{0}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{x})$
for $t=1,\ldots,T$ do
 if $\boldsymbol{c}_{t}\in B_{i}$ then
    $N_{i}(t)=N_{i}(t-1)+1$
    // Increment
      Let $q=\max H(i)$
    // The index of the last event in $B_{i}$
    $\eta_{t}=(4dq_{T})^{-1}\min\left(1,\sqrt{\frac{16(\mu+\beta/\alpha)\log(N_{i}(t)+1)}{N_{i}(t)}}\right)$
    // Update the step size
    $P_{t}=\left(\nabla^{2}\mathcal{R}(\boldsymbol{x}_{q})+\eta_{t}\alpha N_{i}(t)\mathbf{I}_{d}\right)^{-\frac{1}{2}}$
    // Update the perturbation parameter
      Generate $\boldsymbol{\zeta}_{t}$ from $\partial B^{d}$
    // Randomization
    $\boldsymbol{z}_{t}=\boldsymbol{x}_{q}+P_{t}\boldsymbol{\zeta}_{t}$
    // Output
    $y_{t}=f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})+\xi_{t}\quad$
    // Query
    $\boldsymbol{g}_{t}=dy_{t}P_{t}^{-1}\boldsymbol{\zeta}_{t}$
    // Estimate the gradient
    $\boldsymbol{x}_{t}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\Theta}\eta_{t}\sum_{k\in H(i)}\left(\langle\boldsymbol{g}_{k},\boldsymbol{x}\rangle+\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{k-1}\right\rVert^{2}\right)+\mathcal{R}(\boldsymbol{x})$
    // Iterate
    $H(i)=H(i)\cup{t}$
    // Save the index of last event in $B_{i}$
   end if
end for
Algorithm 3
Corollary 9
Let Assumptions 1, 2, and
3 hold. Let $\rho\in[0,1]$ be such that $\mu\leq\texttt{A}d^{\rho}$, where $\texttt{A}>0$ is a numerical constant. Then, the dynamic regret of Algorithm 3 with $K=K(1+\rho/2,1/2,1,\gamma)$ (see (9)), satisfies
$\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\big)\big]\leq C\max\bigg\{L_{1}d^{1+\rho}+L_{2}\sqrt{d^{2+\rho}T},\,$
$\qquad\qquad\qquad\qquad\qquad L_{1}\Big(Lp^{\frac{\gamma}{2}}L_{2}^{-1}\Big)^{\frac{2p}{p+2\gamma}}d^{\frac{2\gamma(1+\rho)-p}{p+2\gamma}}T^{\frac{p}{p+2\gamma}},\,L_{2}^{\frac{2\gamma}{p+2\gamma}}\Big(Lp^{\frac{\gamma}{2}}\Big)^{\frac{p}{p+2\gamma}}d^{\frac{(2+\rho)\gamma}{p+2\gamma}}T^{\frac{p+\gamma}{p+2\gamma}}\bigg\},$
where $L_{1}=\sqrt{\log(T+1)}\log(Td+1)$, $L_{2}=\log(T+1)$, and $C>0$ is a constant depending only on $M,\sigma,\beta/\alpha$.
The main term in the bound of Corollary 9 scales as $T^{\frac{p+\gamma}{p+2\gamma}}$. In Li et al. (2019), this rate was obtained for a different algorithm, under the additional assumption that for each $\boldsymbol{c}\in[0,1]^{p}$ the global minimizer $\operatorname*{arg\,min}_{\boldsymbol{x}\in\mathbb{R}^{d}}f(\boldsymbol{x},\boldsymbol{c})$ is an interior point of the constraint set $\Theta$. While this assumption might be hard to check in practice, it does simplify the analysis considerably. Note also that the multiplicative factors in Corollary 9 exhibit a very mild dependence on both dimensions $p$ and $d$, never exceeding $p^{1/2}d$ in the main term.
5 Lower bound
In this section, along with the classes $\mathcal{F}_{\gamma}(L)$, $\gamma\in(0,1]$, we will consider the class $\mathcal{F}_{0}(L)$, which includes discontinuous functions.
Definition 10
For $L>0$, let $\mathcal{F}_{0}(L)$ be the class of functions such that $f\in\mathcal{F}_{0}(L)$ if
$|f(\boldsymbol{x},\boldsymbol{c})-f(\boldsymbol{x},\boldsymbol{c}^{\prime})|\leq L,\quad\text{for all}\quad\boldsymbol{x}\in\Theta,\,\,\boldsymbol{c},\boldsymbol{c}^{\prime}\in[0,1]^{p}.$ (14)
We will assume that the noises $\{\xi_{t}\}_{t=1}^{T}$ are independent with cumulative distribution function $F$ satisfying the condition
$\int\log\left(\,\mathrm{d}F\left(u\right)/dF\left(u+v\right)\right)\,\mathrm{d}F(u)\leq I_{0}v^{2},\quad\quad|v|<v_{0},$ (15)
for some $0<I_{0}<\infty,0<v_{0}\leq\infty.$ It can be shown that condition (15) is satisfied for $F$ with a sufficiently smooth density and finite Fisher information. If $F$ follows a Gaussian distribution, (15) holds with $v_{0}=\infty$. Note that Gaussian noise satisfies Assumption 3 used to prove the upper bound.
Next, we introduce the distribution of the contexts used in deriving the lower bound. Similar to previous sections, for any $K\in\mathbb{N}$, consider the partition of $0,1]^{p}$, denoted by $\{B_{i}\}_{i=1}^{K^{p}}$, for which we assume that $B_{i}$’s have equal volumes. Moreover, for $i\in[K^{p}]$, let $\mathbf{P}_{i}$ denote the uniform distribution on $G_{i}=(1/2)(B_{i}+\boldsymbol{b}_{i})$, where $\boldsymbol{b}_{i}$ is the barycenter of $B_{i}$. Then, we assume that $\{\boldsymbol{c}_{t}\}_{t=1}^{T}$ are independently distributed according to $\mathbb{P}_{K}$, where
$\mathbb{P}_{K}(\Upsilon)=K^{-p}\sum_{i=1}^{K^{p}}\mathbf{P}_{i}(\Upsilon\cap G_{i}),$ (16)
for $\Upsilon\subseteq[0,1]^{p}$. Distribution $\mathbb{P}_{K}$ represents a challenging scenario for the learning process. Indeed, the contexts are uniformly spread, preventing concentration in any specific region of $0,1]^{p}$. Moreover, they are not clustered near the boundaries of each cell, thus preventing the learner from exploiting proximity between the cells. See Appendix C for further details.
Theorem 11
For $(\alpha,\beta,M,\gamma,L)\in(0,\infty)\times[3\alpha,\infty)\times[\alpha+1,\infty)\times[0,1]\times[0,\infty)$ consider the class of functions $\bar{\mathcal{F}}=\mathcal{F}_{\alpha,\beta}(M)\cap\mathcal{F}_{\gamma}(L)$. Let $\Theta=B^{d}$, and let $\{\boldsymbol{c}_{t}\}_{t=1}^{T}$ be independently distributed according to $\mathbb{P}_{K}$, where $K=\max\left(1,\left\lfloor\left(\min\left(1,L^{2}\right)T\right)^{\frac{1}{p+2\gamma}}\right\rfloor\right)$. Let $\pi\in\Pi$ be any randomized procedure and assume that $\{\boldsymbol{z}_{t}\}_{t=1}^{T}$ are outputs of $\pi$. Then,
$\sup_{f\in\bar{\mathcal{F}}}\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right]\geq\texttt{A}\left(\min\left(1,L^{\frac{2(p+\gamma)}{p+2\gamma}}\right)T^{\frac{p+\gamma}{p+2\gamma}}+\min\left(T,d\sqrt{T}\right)\right),$ (17)
where $\texttt{A}>0$ is a constant that does not depend on $d$, $p$, and $T$.
Applying (17) with $L>0$ and $\gamma=0$ shows that no randomized procedure can achieve sub-linear dynamic regret on the corresponding class $\bar{\mathcal{F}}$. Thus, controlling the increments of $f$ with any $L>0$, in the absence of continuity with respect to $\boldsymbol{c}$, is not sufficient to get a sub-linear dynamic regret.
For $\gamma\in(0,1]$, the bound (17) combined with Corollary 9 implies that Algorithm 3 with $K$ chosen as in Corollary 9 achieves the optimal rate of regret in $T$ up to logarithmic factors. An additional (mild) gap between the upper and lower bounds is due to the discrepancy in the factors depending on $d,p,$ and $L$.
If $L=0$ the bound in (17) scales as $\min(T,d\sqrt{T})$. In this case, the objective functions are constant with respect to the context variable, $f(\boldsymbol{x},\boldsymbol{c}_{1})=f(\boldsymbol{x},\boldsymbol{c}_{t})$, for $\boldsymbol{x}\in\mathbb{R}^{d}$ and $t\in[T]$, so that the contextual regret in (17) is equal to the static regret. Moreover, by convexity, both regrets are bounded from below by the optimization error corresponding to $f(\cdot,\boldsymbol{c}_{1})$ multiplied by $T$. Using this remark we notice that, for $L=0$, the bound (17) agrees with the lower bound on the optimization error for strongly convex and smooth functions, known to be of the order $d/\sqrt{T}$, see Shamir (2013); Akhavan et al. (2020, 2024).
Furthermore, note that since the class of strongly convex functions is included in the class of convex functions, (17) is a valid lower bound for convex contextual bandits. Together with Corollary 4, this implies that the minimax optimal rate as a function of $T$ for convex contextual bandits is of the order $T^{\frac{p+\gamma}{p+2\gamma}}$ up to logarithmic factors.
Theorem 11 can be stated in a more general manner. In the definition of $\Pi$, the randomized procedures are restricted to query only within the constraint set $\Theta$. However, in the proof of the lower bound, we did not impose such a restriction, allowing $\pi$ to query anywhere within $\mathbb{R}^{d}$. Thus, Theorem 11 remains valid for this broader set of algorithms. Furthermore, the proof of Theorem 11 can be extended to hold for any convex body $\Theta$ and not only for $\Theta=B^{d}$. Indeed, by shifting and scaling the unit Euclidean ball, we can embed it in $\Theta$. Subsequently, we can apply the same shifting and scaling to the action argument of the set of functions $f_{\boldsymbol{\omega},\boldsymbol{\tau}}$ in the proof of Theorem 11.
Disclosure of Funding
The research of Arya Akhavan was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number EP/Y028333/1].
Appendix
In Appendix A, we present the proof of Theorem 2. Appendices B and C begin with a set of auxiliary and preliminary lemmas, followed by the proofs of the main results: Theorem 8, Corollary 9, and Theorem 11. In Appendix D, we collect some properties of self-concordant barrier functions.
Appendix A Proof of Theorem 2
Proof [Proof of Theorem 2] For $i\in[K^{p}]$, let $N_{i}(T)$ the number of $\boldsymbol{c}_{t}$’s belonging to $B_{i}$. Denote $i_{1},\dots,i_{N_{i}(T)}$ as the set of all indices such that $\boldsymbol{c}_{i_{j}}\in B_{i}$ for all $j\in[N_{i}(T)]$. We have
$\sum_{t=1}^{T}\left(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right)=\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})\right).$
For $i\in[K^{p}]$, let $\boldsymbol{b}_{i}\in B_{i}$ be the barycenter of the cube $B_{i}$, and $\boldsymbol{x}^{*}_{i}\in\operatorname*{arg\,min}_{\boldsymbol{x}\in\Theta}f(\boldsymbol{x},\boldsymbol{b}_{i})$. Since (7) holds it follows that
$\mathbf{E}\left[\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}_{i}^{*},\boldsymbol{c}_{i_{j}})\right)|N_{i}(T)\right]\leq F(N_{i}(T)){F_{1}(N_{i}(T))\leq F(N_{i}(T))F_{1}(T)},$
where $F$ is a concave function and $F_{1}$ is a non-decreasing function. Note that $\mathbf{E}\left[N_{i}(T)\right]=p_{i}T$, for some $p_{i}\in[0,1]$, such that $\sum_{i=1}^{K^{p}}p_{i}=1$. By taking expectations from both sides and applying Jensen’s inequality we get
$\mathbf{E}\left[\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})\right)\right]$ $\leq F(p_{i}T){F_{1}(T)}.$
Summing both sides over $i\in[K^{p}]$ and using Jensen’s inequality again we obtain
$\mathbf{E}\left[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})\right)\right]$ $\leq\sum_{i=1}^{K^{p}}F(p_{i}T){F_{1}(T)}$
$\leq K^{p}F\left(\sum_{i=1}^{K^{p}}\frac{p_{i}T}{K^{p}}\right){F_{1}(T)}$
$=K^{p}F\left(\frac{T}{K^{p}}\right){F_{1}(T)}.$ (18)
If, in addition, $\pi\in\Pi(\mathcal{F},\tau_{1},\tau_{2},\tau_{3},\mathrm{T}_{0})$, then for $T\geq K^{p}\,\mathrm{T}_{0}$ we get that
$\mathbf{E}\left[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})\right)\right]\leq Cd^{\tau_{1}}T^{\tau_{2}}\log^{\tau_{3}}(T+1)K^{p(1-\tau_{2})}.$ (19)
Let $\boldsymbol{z}^{*}_{i_{j}}\in\operatorname*{arg\,min}_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{i_{j}})$. We can write
$\mathbf{E}\left[\sum_{t=1}^{T}\left(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right)\right]$ $=\mathbf{E}\left[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})+f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{z}^{*}_{i_{j}},\boldsymbol{c}_{i_{j}})\right)\right]$
$=\mathbf{E}\bigg[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\bigg(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})+f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}_{i}^{*},\boldsymbol{b}_{i})$
$\quad+f(\boldsymbol{x}_{i}^{*},\boldsymbol{b}_{i})-f(\boldsymbol{z}^{*}_{i_{j}},\boldsymbol{c}_{i_{j}})\bigg)\bigg].$
Since $f(\boldsymbol{x}_{i}^{*},\boldsymbol{b}_{i})\leq f(\boldsymbol{z}_{i_{j}}^{*},\boldsymbol{b}_{i})$, we have
$\mathbf{E}\left[\sum_{t=1}^{T}\left(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right)\right]$ $\leq\mathbf{E}\bigg[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\bigg(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})+f(\boldsymbol{x}_{i}^{*},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}_{i}^{*},\boldsymbol{b}_{i})$
$\quad+f(\boldsymbol{z}^{*}_{i_{j}},\boldsymbol{b}_{i})-f(\boldsymbol{z}^{*}_{i_{j}},\boldsymbol{c}_{i_{j}})\bigg)\bigg]$
$\leq\mathbf{E}\left[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})+2L\left\lVert\boldsymbol{c}_{i_{j}}-\boldsymbol{b}_{i}\right\rVert^{\gamma}\right)\right],$
where the last inequality is due to Assumption 1. Let $\delta=\max_{i\in[K^{p}]}\max_{\boldsymbol{x},\boldsymbol{y}\in B_{i}}\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert$. Since all $B_{i}$’s are hypercubes with edges of length $1/K$ we have $\delta=\sqrt{p}/K$, so that
$\mathbf{E}\left[\sum_{t=1}^{T}\left(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right)\right]$ $\leq\mathbf{E}\left[\sum_{i=1}^{K^{p}}\sum_{j=1}^{N_{i}(T)}\left(f(\boldsymbol{z}_{i_{j}},\boldsymbol{c}_{i_{j}})-f(\boldsymbol{x}^{*}_{i},\boldsymbol{c}_{i_{j}})\right)\right]+2TL\left(\frac{\sqrt{p}}{K}\right)^{\gamma}.$ (20)
Combining (18) and (20), gives
$\mathbf{E}\left[\sum_{t=1}^{T}\left(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right)\right]$ $\leq K^{p}F\left(\frac{T}{K^{p}}\right)+2TL\left(\frac{\sqrt{p}}{K}\right)^{\gamma},$
which concludes the proof of (10). In order to prove the bound (2), we note that, by (19) and (20),
$\mathbf{E}\left[\sum_{t=1}^{T}\left(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\right)\right]$ $\leq Cd^{\tau_{1}}T^{\tau_{2}}\log^{\tau_{3}}(T+1)K^{p(1-\tau_{2})}+2TL\left(\frac{\sqrt{p}}{K}\right)^{\gamma}.$
The two terms on the right hand side are of the same order of magnitude if $K=K(\tau_{1},\tau_{2},\tau_{3},\gamma)$. Under this choice of $K$ we get
$\mathbf{E}\bigg[\sum_{t=1}^{T}\bigg(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}$ $f(\boldsymbol{z},\boldsymbol{c}_{t})\bigg)\bigg]\leq C^{\prime}\max\bigg(d^{\tau_{1}}T^{\tau_{2}}\log^{\tau_{3}}(T+1),$
$\left(Lp^{\frac{\gamma}{2}}\right)^{\frac{p(1-\tau_{2})}{p(1-\tau_{2})+\gamma}}\left(d^{\tau_{1}}\log^{\tau_{3}}(T+1)\right)^{\frac{\gamma}{p(1-\tau_{2})+\gamma}}T^{\frac{(p-\gamma)(1-\tau_{2})+\gamma}{p(1-\tau_{2})+\gamma}}\bigg),$
where $C^{\prime}>0$ is a constant that does not depend on $T,p$ and $d$.
Appendix B Proof of Theorem 8 and Corollary 9
Lemma 12
Let $\{\eta_{t}\}_{t=1}^{T}$ be as in Algorithm 2. Then, for all $t\geq 2$ we have
$1-\frac{\eta_{t}}{\eta_{t-1}}\leq\mathbf{1}\left(t\geq 16\mu\right)\sqrt{\frac{1}{t}}.$ (21)
Proof Note that $t/\log(t+1)$ is an increasing function for all $t\geq 1$. Let $t_{0}$ be the smallest positive integer such that $t_{0}/\log(t_{0}+1)\geq 16(\mu+\beta/\alpha)$, and note that
$\eta_{t}=(4dq_{T})^{-1},\quad\text{for }t<t_{0},$
and
$\eta_{t}=(4dq_{T})^{-1}\left(\frac{16(\mu+\beta/\alpha)\log(t+1)}{t}\right)^{\frac{1}{2}},\quad\text{for }t\geq t_{0},$
where $q_{T}=M+2\sigma\sqrt{\log(T+1)}$. If $t<t_{0}$ then $1-\eta_{t}/\eta_{t-1}=0$. For the case $t=t_{0}$, we can write
$1-\frac{\eta_{t}}{\eta_{t-1}}=\frac{\sqrt{t_{0}}-\sqrt{16(\mu+\beta/\alpha)\log(t_{0}+1)}}{\sqrt{t_{0}}}.$
Note that $t_{0}\leq 16\left(\mu+\beta/\alpha\right)\log(t_{0}+1)+1$. Thus,
$1-\frac{\eta_{t}}{\eta_{t-1}}\leq\frac{\sqrt{t_{0}}-\sqrt{t_{0}-1}}{\sqrt{t_{0}}}\leq\sqrt{\frac{1}{t_{0}}}=\sqrt{\frac{1}{t}}.$
For $t>t_{0}$ we have
$1-\frac{\eta_{t}}{\eta_{t-1}}=1-\sqrt{\frac{t-1}{t}}\sqrt{\frac{\log(t+1)}{\log(t)}}\leq\frac{\sqrt{t}-\sqrt{t-1}}{\sqrt{t}}\leq\sqrt{\frac{1}{t}}.$
By combing the above inequalities we get
$1-\frac{\eta_{t}}{\eta_{t-1}}\leq\mathbf{1}\left(t\geq t_{0}\right)\sqrt{\frac{1}{t}}.$
Since $16(\mu+\beta/\alpha)\geq 16$, we have $t_{0}\geq 2$, and consequently $\log(t_{0}+1)\geq 1$. Thus,
$t_{0}\geq 16(\mu+\frac{\beta}{\alpha})\log(t_{0}+1)\geq 16\mu.$
Therefore, for all $t\in[T]$ we have
$\mathbf{1}\left(t\geq t_{0}\right)\leq\mathbf{1}\left(t\geq 16\mu\right),$
which concludes the proof.
Before proceeding with the proofs, let us recall the concept of the Newton decrement of a self-concordant function. Let $\Psi:\text{int}(\Theta)\to\mathbb{R}$ be a self-concordant function. The Newton decrement of $\Psi$ at $\boldsymbol{x}\in\text{int}(\Theta)$ is defined as
$\lambda(\Psi,\boldsymbol{x})=\max_{\boldsymbol{h}\in\mathbb{R}^{d}}\{\nabla\Psi(\boldsymbol{x})[\boldsymbol{h}]|\nabla^{2}\Psi(\boldsymbol{x})[\boldsymbol{h},\boldsymbol{h}]\leq 1\}.$ (22)
In the case that $\nabla^{2}\Psi(\boldsymbol{x})$ is positive definite, for all $\boldsymbol{x}\in\text{int}(\Theta)$ (22) can be written as
$\lambda(\Psi,\boldsymbol{x})=\sqrt{(\nabla\Psi(\boldsymbol{x}))^{\top}(\nabla^{2}\Psi(\boldsymbol{x}))^{-1}\nabla\Psi(\boldsymbol{x})}.$ (23)
The proof of the equivalency of (22) and (23) can be found in (Nemirovski, 2004, Chapter 2, Section IVa). Moreover, if $\Psi$ is a $\mu$-self-concordant barrier then $\lambda(\Psi,\boldsymbol{x})\leq\mu^{1/2}$ for all $\boldsymbol{x}\in\text{int}(\Theta)$.
The next lemma plays a key role in our analysis.
Lemma 13
Let $\Psi:\text{int}(\Theta)\to\mathbb{R}$ be a self-concordant function. Then $\Psi$ attains its minimum on $\text{int}(\Theta)$ if and only if there exists $\boldsymbol{x}\in\text{int}(\Theta)$ with $\lambda(\Psi,\boldsymbol{x})<1$. Moreover, if $\lambda(\Psi,\boldsymbol{x})<1$, then
$((\boldsymbol{x}-\boldsymbol{x}^{*})^{\top}\nabla^{2}\Psi(\boldsymbol{x})(\boldsymbol{x}-\boldsymbol{x}^{*}))^{\frac{1}{2}}\leq\frac{\lambda(\Psi,\boldsymbol{x})}{1-\lambda(\Psi,\boldsymbol{x})},$
where $\boldsymbol{x}^{*}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\text{int}(\Theta)}\Psi(\boldsymbol{x})$.
For a proof of Lemma 13, we refer the reader to (Nemirovski, 2004, Chapter 2, Section VIII).
Lemma 14
Define the event
$\Lambda_{T}=\{|\xi_{t}|\leq 2\sigma\sqrt{\log(T+1)}\,\text{ for all }t\in[T]\}.$
Let $\{\boldsymbol{x}_{t}\}_{t=0}^{T}$ be the iterates defined in Algorithm 2. If $\Lambda_{T}$ holds, for all $t\in[T]$ we have
$\left\lVert\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\right\rVert_{t}\leq 2\left(\eta_{t}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}+\sqrt{\frac{\mu}{t}}\right),$ (24)
where we use the notation $\left\lVert\boldsymbol{u}\right\rVert_{t}:=\sqrt{\boldsymbol{u}^{\top}P_{t}^{-2}\boldsymbol{u}}$ and $\left\lVert\boldsymbol{u}\right\rVert_{t}^{*}:=\sqrt{\boldsymbol{u}^{\top}P_{t}^{2}\boldsymbol{u}}$   for all $\boldsymbol{u}\in\mathbb{R}^{d}$.
Proof Let $\tilde{\Phi}_{t}:\text{int}(\Theta)\to\mathbb{R}$ be such that
$\tilde{\Phi}_{t}(\boldsymbol{u})=\eta_{t}\left(\sum_{k=1}^{t}\langle\boldsymbol{g}_{k},\boldsymbol{u}\rangle+\frac{\alpha}{2}\sum_{k=1}^{t}\left\lVert\boldsymbol{u}-\boldsymbol{x}_{k-1}\right\rVert^{2}\right)+\mathcal{R}(\boldsymbol{u}).$
Note that $\tilde{\Phi}_{t}$ is a self-concordant function with $\boldsymbol{x}_{t}\in\operatorname*{arg\,min}_{\boldsymbol{x}\in\text{int}(\Theta)}\tilde{\Phi}_{t}(\boldsymbol{x})$, and $\nabla^{2}\tilde{\Phi}_{t}(\boldsymbol{x}_{t-1})=P_{t}^{-2}$. Consider the Newton decrement $\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1})$, cf. (23). The first part of the proof is dedicated to providing an upper bound for $\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1})$. Then we use Lemma 13 to prove (24). For $t\in[T]$ we have
$\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1})=\sqrt{(\nabla\tilde{\Phi}_{t}(\boldsymbol{x}_{t-1}))^{\top}(\nabla^{2}\tilde{\Phi}_{t}(\boldsymbol{x}_{t-1}))^{-1}\nabla\tilde{\Phi}_{t}(\boldsymbol{x}_{t-1})}=\left\lVert\nabla\tilde{\Phi}_{t}(\boldsymbol{x}_{t-1})\right\rVert_{t}^{*}.$ (25)
For $t=1$, we can write
$\left\lVert\nabla\tilde{\Phi}_{1}(\boldsymbol{x}_{0})\right\rVert_{1}^{*}=\left\lVert\eta_{1}\langle\boldsymbol{g}_{1},\boldsymbol{x}_{0}\rangle+\nabla\mathcal{R}(\boldsymbol{x}_{0})\right\rVert\leq\eta_{1}\left\lVert\boldsymbol{g}\right\rVert_{1}^{*},$ (26)
where the last inequality follows from the facts that $\boldsymbol{x}_{0}=\operatorname*{arg\,min}_{\boldsymbol{x}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{x})$ and $\mathcal{R}(\boldsymbol{x}_{0})=0$.
Moreover, for $t\geq 2$ we have
$\nabla\tilde{\Phi}_{t}(\boldsymbol{x}_{t-1})$ $=\frac{\eta_{t}}{\eta_{t-1}}\left(\eta_{t-1}\sum_{k=1}^{t-1}\boldsymbol{g}_{k}+\eta_{t-1}\alpha\sum_{k=1}^{t-1}(\boldsymbol{x}-\boldsymbol{x}_{k-1})+\nabla\mathcal{R}(\boldsymbol{x}_{t-1})\right)+\eta_{t}\boldsymbol{g}_{t}$
$\phantom{00000}+\left(1-\frac{\eta_{t}}{\eta_{t-1}}\right)\nabla\mathcal{R}(\boldsymbol{x}_{t-1})$
$=\frac{\eta_{t}}{\eta_{t-1}}\nabla\tilde{\Phi}_{t-1}(\boldsymbol{x}_{t-1})+\eta_{t}\boldsymbol{g}_{t}+\left(1-\frac{\eta_{t}}{\eta_{t-1}}\right)\nabla\mathcal{R}(\boldsymbol{x}_{t-1})$
$=\eta_{t}\boldsymbol{g}_{t}+\left(1-\frac{\eta_{t}}{\eta_{t-1}}\right)\nabla\mathcal{R}(\boldsymbol{x}_{t-1}).$ (27)
Using (25) – (26) we find that, for all $t\in[T]$,
$\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1})=\left\lVert\nabla\tilde{\Phi}_{t}(\boldsymbol{x}_{t-1})\right\rVert_{t}^{*}\leq\eta_{t}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}+\left(1-\frac{\eta_{t}}{\eta_{t-1}}\right)\left\lVert\nabla\mathcal{R}(\boldsymbol{x}_{t-1})\right\rVert_{t}^{*}.$
Thanks to Lemma 12 we can bound the term $1-\eta_{t}/\eta_{t-1}$ to obtain
$\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1})\leq\eta_{t}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}+\left\lVert\nabla\mathcal{R}(\boldsymbol{x}_{t-1})\right\rVert_{t}^{*}\mathbf{1}\left(t\geq 16\mu\right)\sqrt{\frac{1}{t}}.$
On the other hand, since $\mathcal{R}$ is a $\mu$-self-concordant function we have
$(\left\lVert\nabla\mathcal{R}(\boldsymbol{x}_{t-1})\right\rVert_{t}^{*})^{2}\leq\nabla\mathcal{R}(\boldsymbol{x}_{t-1})^{\top}(\nabla^{2}\mathcal{R}(\boldsymbol{x}_{t-1}))^{-1}\nabla\mathcal{R}(\boldsymbol{x}_{t-1})=\lambda^{2}(\mathcal{R},\boldsymbol{x}_{t-1})\leq\mu,$
so that
$\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1})\leq\eta_{t}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}+\mathbf{1}\left(t\geq 16\mu\right)\sqrt{\frac{\mu}{t}}\leq\eta_{t}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}+\frac{1}{4}.$ (28)
Since we assumed that $\Lambda_{T}$ holds, we can write
$\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}=\left\lVert d(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})+\xi_{t})P_{t}^{-1}\boldsymbol{\zeta}_{t}\right\rVert_{t}^{*}\leq d(|f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})|+|\xi_{t}|)\leq d(M+2\sigma\sqrt{\log(T+1)}).$ (29)
By the definition of $\eta_{t}$, for all $t\in[T]$ we have that $\eta_{t}\leq(4d(M+2\sigma\sqrt{\log(T+1)}))^{-1}$ which implies that $\eta_{t}\left\lVert\boldsymbol{g}\right\rVert_{t}^{*}\leq 1/4$. Using this inequality in (28) gives $\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1})\leq 1/2$. Therefore, by Lemma 13 we can write
$\left\lVert\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\right\rVert_{t}=\left\lVert\boldsymbol{x}_{t-1}-\operatorname*{arg\,min}_{\boldsymbol{x}\in\text{int}(\Theta)}\tilde{\Phi}_{t}(\boldsymbol{x})\right\rVert_{t}\leq 2\lambda(\tilde{\Phi}_{t},\boldsymbol{x}_{t-1}).$
The last display together with (28) yields
$\left\lVert\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\right\rVert$ $\leq 2\left(\eta_{t}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}+\mathbf{1}\left(t\geq 16\mu\right)\sqrt{\frac{\mu}{t}}\right)\leq 2\left(\eta_{t}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}+\sqrt{\frac{\mu}{t}}\right).$
Lemma 15
Consider a function $h:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$, and set $h_{t}(\boldsymbol{x})=h(\boldsymbol{x},\boldsymbol{c}_{t})$ for $t\geq 1$. Assume that $\{h_{t}\}_{t=1}^{T}$ is a sequence of $\alpha$-strongly convex functions for some $\alpha>0$. Let $\{\boldsymbol{x}_{t}\}_{t=0}^{T}$ be the iterates defined in Algorithm 2. Let $\boldsymbol{g}_{t}$’s be random vectors in $\mathbb{R}^{d}$ such that $\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda_{T}\right]=\nabla h_{t}(\boldsymbol{x}_{t-1})$ for $t\geq 1$. Then, for any $\boldsymbol{x}\in\text{int}(\Theta)$ we have
$\sum_{t=1}^{T}\mathbf{E}\left[h_{t}(\boldsymbol{x}_{t-1})-h_{t}(\boldsymbol{x})|\Lambda_{T}\right]\leq 2dq_{T}\left(dq_{T}\sum_{t=1}^{T}\eta_{t}+2\sqrt{\mu T}\right)+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x}),$
where $q_{T}=M+2\sigma\sqrt{\log(T+1)}$.
Proof Fix $\boldsymbol{x}\in\text{int}(\Theta)$. Since $h_{t}$ is a strongly convex function, for any $\boldsymbol{x}\in\Theta$ we can write
$\sum_{t=1}^{T}\left(h_{t}(\boldsymbol{x}_{t-1})-h_{t}(\boldsymbol{x})\right)\leq\sum_{t=1}^{T}\langle\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda_{T}\right],\boldsymbol{x}_{t-1}-\boldsymbol{x}\rangle-\frac{\alpha}{2}\sum_{t=1}^{T}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{t-1}\right\rVert^{2}.$ (30)
To simplify the notation, we rewrite $\boldsymbol{x}_{t}$ as follows
$\boldsymbol{x}_{t}$ $=\operatorname*{arg\,min}_{\boldsymbol{u}\in\Theta}\sum_{k=1}^{t}\langle\boldsymbol{g}_{k},\boldsymbol{u}\rangle+\underbrace{\frac{\alpha}{2}\sum_{k=1}^{t}\left\lVert\boldsymbol{u}-\boldsymbol{x}_{k-1}\right\rVert^{2}+\eta_{t}^{-1}\mathcal{R}(\boldsymbol{u})}_{=\vcentcolon R_{t}(\boldsymbol{u})}$
$=\operatorname*{arg\,min}_{\boldsymbol{u}\in\Theta}\underbrace{\sum_{k=1}^{t}\langle\boldsymbol{g}_{k},\boldsymbol{u}\rangle+R_{t}(\boldsymbol{u})}_{=\vcentcolon\Phi_{t}(\boldsymbol{u})}=\operatorname*{arg\,min}_{\boldsymbol{u}\in\Theta}{\Phi_{t}(\boldsymbol{u})}.$
Moreover, set $\Phi_{0}(\boldsymbol{u})=\mathcal{R}(\boldsymbol{u})$ for all $\boldsymbol{u}\in\text{int}(\Theta)$. Since $\boldsymbol{x}_{0}=\operatorname*{arg\,min}_{\boldsymbol{u}\in\text{int}(\Theta)}\Phi_{0}(\boldsymbol{u})$ and we assume that $\mathcal{R}(\boldsymbol{x}_{0})=0$, we have
$-\sum_{t=1}^{T}\langle\boldsymbol{g}_{t},\boldsymbol{x}\rangle$ $=-\Phi_{T}(\boldsymbol{x})+\Phi_{T}(\boldsymbol{x}_{T})-\Phi_{T}(\boldsymbol{x}_{T})+R_{T}(\boldsymbol{x})$
$\leq\sum_{t=1}^{T}\left(\Phi_{t-1}(\boldsymbol{x}_{t-1})-\Phi_{t}(\boldsymbol{x}_{t})\right)+R_{T}(\boldsymbol{x}),$
where the last display is due to the fact that $\Phi_{T}(\boldsymbol{x}_{T})\leq\Phi_{T}(\boldsymbol{x})$. Subsequently,
$\sum_{t=1}^{T}\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}-\boldsymbol{x}\rangle\leq\sum_{t=1}^{T}\left(\underbrace{\Phi_{t-1}(\boldsymbol{x}_{t-1})-\Phi_{t}(\boldsymbol{x}_{t})+\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}\rangle}_{\text{term I}}\right)+R_{T}(\boldsymbol{x}).$ (31)
In order to control term I, we first derive an alternative expression for $\Phi_{t}$. By the definition of $\Phi_{t}$,
$\Phi_{t}(\boldsymbol{x}_{t})=\sum_{k=1}^{t}\langle\boldsymbol{g}_{k},\boldsymbol{x}_{t}\rangle+R_{t}(\boldsymbol{x}_{t})$ $=\sum_{k=1}^{t-1}\langle\boldsymbol{g}_{k},\boldsymbol{x}_{t}\rangle+\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}\rangle+R_{t-1}(\boldsymbol{x}_{t})+\left(R_{t}(\boldsymbol{x}_{t})-R_{t-1}(\boldsymbol{x}_{t})\right)$
$=\Phi_{t-1}(\boldsymbol{x}_{t})+\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}\rangle+\left(R_{t}(\boldsymbol{x}_{t})-R_{t-1}(\boldsymbol{x}_{t})\right),$
so that
term I $=\left(\Phi_{t-1}(\boldsymbol{x}_{t-1})+\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}\rangle\right)-\left(\Phi_{t-1}(\boldsymbol{x}_{t})+\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}\rangle\right)+R_{t-1}(\boldsymbol{x}_{t})-R_{t}(\boldsymbol{x}_{t})$
$\leq\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\rangle,$ (32)
where the last inequality is due to the fact that $R_{t-1}(\boldsymbol{x}_{t})\leq R_{t}(\boldsymbol{x}_{t})$, and $\Phi_{t-1}(\boldsymbol{x}_{t-1})=\min_{\boldsymbol{z}\in\Theta}\Phi_{t-1}(\boldsymbol{z})\leq\Phi_{t-1}(\boldsymbol{x}_{t})$. From (31) and (32) we obtain
$\sum_{t=1}^{T}\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}-\boldsymbol{x}\rangle\leq\sum_{t=1}^{T}\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\rangle+\frac{\alpha}{2}\sum_{t=1}^{T}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{t-1}\right\rVert^{2}+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x}).$
Taking expectations conditioned on the event $\Lambda_{T}$ gives
$\mathbf{E}\left[\sum_{t=1}^{T}\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}-\boldsymbol{x}\rangle|\Lambda_{T}\right]$ $\leq\mathbf{E}\left[\sum_{t=1}^{T}\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\rangle+\frac{\alpha}{2}\sum_{t=1}^{T}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{t-1}\right\rVert^{2}|\Lambda_{T}\right]+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x})$
$\leq\mathbf{E}\left[\sum_{t=1}^{T}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}\left\lVert\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\right\rVert_{t}+\frac{\alpha}{2}\sum_{t=1}^{T}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{t-1}\right\rVert^{2}|\Lambda_{T}\right]+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x}).$
Furthermore, by the tower rule,
$\mathbf{E}\left[\sum_{t=1}^{T}\langle\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda_{T}\right],\boldsymbol{x}_{t-1}-\boldsymbol{x}\rangle\big|\Lambda_{T}\right]$ $\leq\mathbf{E}\left[\sum_{t=1}^{T}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}\left\lVert\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\right\rVert_{t}+\frac{\alpha}{2}\sum_{t=1}^{T}\left\lVert\boldsymbol{x}-\boldsymbol{x}_{t-1}\right\rVert^{2}|\Lambda_{T}\right]$
$\phantom{0000000}+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x}).$
Using the assumption that $\mathbf{E}\left[\boldsymbol{g}_{t}|\,\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda_{T}\right]=\nabla h_{t}(\boldsymbol{x}_{t-1})$, the fact that $\{h_{t}\}_{t=1}^{T}$ are $\alpha$-strongly convex functions, and (30), yields
$\mathbf{E}\left[\sum_{t=1}^{T}(h_{t}(\boldsymbol{x}_{t-1})-h_{t}(\boldsymbol{x}))|\Lambda_{T}\right]\leq\mathbf{E}\left[\sum_{t=1}^{T}\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}\left\lVert\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\right\rVert_{t}|\Lambda_{T}\right]+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x}).$ (33)
We now apply Lemma 14 to bound the value $\left\lVert\boldsymbol{x}_{t-1}-\boldsymbol{x}_{t}\right\rVert_{t}$ on the event $\Lambda_{T}$. It follows that
$\mathbf{E}\left[\sum_{t=1}^{T}(h_{t}(\boldsymbol{x}_{t-1})-h_{t}(\boldsymbol{x}))|\Lambda_{T}\right]\leq 2\left(\sum_{t=1}^{T}\eta_{t}\mathbf{E}\left[(\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*})^{2}|\Lambda_{T}\right]+\sqrt{\frac{\mu}{t}}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}|\Lambda_{T}\right]\right)+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x}).$ (34)
Recall that $q_{T}=M+2\sigma\sqrt{\log(T+1})$, and under $\Lambda_{T}$, the uniform bound (29) gives $\left\lVert\boldsymbol{g}_{t}\right\rVert_{t}^{*}\leq dq_{T}$. Hence,
$\sum_{t=1}^{T}\mathbf{E}\left[h_{t}(\boldsymbol{x}_{t-1})-h_{t}(\boldsymbol{x})|\Lambda_{T}\right]$ $\leq 2\left(d^{2}q_{T}^{2}\sum_{t=1}^{T}\eta_{t}+dq_{T}\sum_{t=1}^{T}\sqrt{\frac{\mu}{t}}\right)+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x})$
$\leq 2dq_{T}\left(dq_{T}\sum_{t=1}^{T}\eta_{t}+2\sqrt{\mu T}\right)+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{x}).$
Lemma 16
(Hazan and Levy, 2014, Corollary 6) Let $A\in\mathbb{R}^{d\times d}$ be an invertible matrix. Let $\boldsymbol{U}$ be distributed uniformly on $B^{d}$. Let ${\sf f}_{A}(\boldsymbol{x},\boldsymbol{c})\triangleq\mathbf{E}\left[f(\boldsymbol{x}+A\boldsymbol{U},\boldsymbol{c})\right]$ for $\boldsymbol{x}\in\mathbb{R}^{d}$ and $\boldsymbol{c}\in[0,1]^{p}$. Then,
$\nabla_{\boldsymbol{x}}{\sf f}_{A}(\boldsymbol{x},\boldsymbol{c})=\mathbf{E}\left[df(\boldsymbol{x}+A\boldsymbol{\zeta},\boldsymbol{c})A^{-1}\boldsymbol{\zeta}\right],$
where $\boldsymbol{\zeta}$ is uniformly distributed on $\partial B^{d}$. If Assumption 2(i) holds, then ${\sf f}_{A}(\cdot,\boldsymbol{c})$ is $\alpha$-strongly convex on $\mathbb{R}^{d}$ for all $\boldsymbol{c}\in[0,1]^{p}$.
Lemma 17
Let Assumption 3(ii) hold. Let $\boldsymbol{U}$ be distributed uniformly on $B^{d}$. For $t\geq 1$, let ${\sf f}_{t,P}:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$ be a surrogate function defined by the relation
${\sf f}_{t,P}(\boldsymbol{x},\boldsymbol{c})=\mathbf{E}\left[f(\boldsymbol{x}+P_{t}\boldsymbol{U},\boldsymbol{c})|\boldsymbol{x}_{t-1}\right].$ (35)
Let $\{\boldsymbol{g}_{t}\}_{t=1}^{T}$ and $\{\boldsymbol{x}_{t}\}_{t=0}^{T-1}$ be defined as in Algorithm 2, and $\Lambda$ be any event that only depends on $\{\xi_{t}\}_{t=1}^{T}$. Then for all $t\in[T]$ we have
$\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda\right]=\nabla_{\boldsymbol{x}}{\sf f}_{t,P}(\boldsymbol{x}_{t-1},\boldsymbol{c}_{t}).$
Proof By the definition of $\boldsymbol{g}_{t}$ we have
$\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda\right]$ $=d\mathbf{E}\left[\left(f(\boldsymbol{x}_{t-1}+P_{t}\boldsymbol{\zeta}_{t},\boldsymbol{c}_{t})+\xi_{t}\right)P_{t}^{-1}\boldsymbol{\zeta}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda\right]$
$=$ $d\mathbf{E}\left[f(\boldsymbol{x}_{t-1}+P_{t}\boldsymbol{\zeta}_{t},\boldsymbol{c}_{t})P_{t}^{-1}\boldsymbol{\zeta}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t}\right]+\underbrace{d\mathbf{E}[\xi_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda]P_{t}^{-1}\mathbf{E}\left[\boldsymbol{\zeta}_{t}\right]}_{=0},$
where the last equality is due to Assumption 3(ii). Now, by Lemma 16 we have
$d\mathbf{E}\left[f(\boldsymbol{x}_{t-1}+P_{t}\boldsymbol{\zeta}_{t},\boldsymbol{c}_{t})P_{t}^{-1}\boldsymbol{\zeta}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t}\right]=\nabla_{\boldsymbol{x}}{\sf f}_{t,P}(\boldsymbol{x}_{t-1},\boldsymbol{c}_{t}),$
which concludes the proof.
Lemma 18
(Hazan and Levy, 2014, Lemma 4) For $\mu>0$ let $\mathcal{R}:\text{int}(\Theta)\to\mathcal{R}$ be a $\mu$-self-concordant barrier with $\min_{\boldsymbol{x}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{x})=0$. Then, for all $\boldsymbol{x},\boldsymbol{y}\in\text{int}(\Theta)$ we have
$\mathcal{R}(\boldsymbol{x})-\mathcal{R}(\boldsymbol{y})\leq\mu\log\left(\frac{1}{1-\pi_{\boldsymbol{x}}(\boldsymbol{y})}\right),$
where $\pi_{\boldsymbol{x}}(\boldsymbol{y})=\inf\{t\geq 0:\boldsymbol{x}+t^{-1}(\boldsymbol{y}-\boldsymbol{x})\}$ is the Minkowsky function. Particularly, for $\boldsymbol{x}\in\text{int}(\Theta)$ if $\boldsymbol{x}_{0}\in\operatorname*{arg\,min}_{\boldsymbol{u}\in\text{int}(\Theta)}\mathcal{R}(\boldsymbol{u})$, and $\boldsymbol{x}^{\prime}=(1-1/T)\boldsymbol{x}+(1/T)\boldsymbol{x}_{0}$ then
$\mathcal{R}(\boldsymbol{x}^{\prime})\leq\mu\log\left(T\right).$ (36)
Proof [Proof of Theorem 8] Let $\boldsymbol{z}^{\prime}=(1-1/T)\boldsymbol{z}+(1/T)\boldsymbol{x}_{0}$. Note that
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z},\boldsymbol{c}_{t})\right]$ $=\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})\right]+\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})-f(\boldsymbol{z},\boldsymbol{c}_{t})\right]$
$\leq\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})\right]+\frac{1}{T}\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{x}_{0},\boldsymbol{c}_{t})-f(\boldsymbol{z},\boldsymbol{c}_{t})\right]$ (37)
$\leq\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})\right]+2M,$ (38)
where (37) is due to convexity of $f(\cdot,\boldsymbol{c})$. Invoking the event $\Lambda_{T}$ we can write
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})\right]$ $=\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}^{c}\right]\mathbf{P}[\Lambda_{T}^{c}]+\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]\mathbf{P}[\Lambda_{T}].$
Since $\{\xi_{t}\}_{t=1}^{T}$ are $\sigma$-sub-Gaussian we have
$\mathbf{P}\left[\Lambda_{T}^{c}\right]\leq\sum_{t=1}^{T}\mathbf{P}\left[|\xi_{t}|>2\sigma\sqrt{\log(T+1})\right]\leq 2\sum_{t=2}^{T+1}T^{-2}=2T^{-1}.$
Thus,
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})\right]\leq 4M+\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right].$ (39)
Next, introducing the surrogate function ${\sf f}_{t,P}(\boldsymbol{x},\boldsymbol{c})=\mathbf{E}\left[f(\boldsymbol{x}+P_{t}\boldsymbol{U},\boldsymbol{c})|\boldsymbol{x}_{t-1}\right]$ we consider the decomposition
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]$
[…]
$\hskip 56.9055pt+\underbrace{\sum_{t=1}^{T}\mathbf{E}\left[{\sf f}_{t,P}(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]}_{\text{term III}}+\underbrace{\sum_{t=1}^{T}\mathbf{E}\left[{\sf f}_{t,P}(\boldsymbol{x}_{t-1},\boldsymbol{c}_{t})-{\sf f}_{t,P}(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]}_{\text{term IV}}.$
In what follows, we bound each term in this decomosition separately. For term I, by Assumption 2(ii) we have
$\text{term I}=\mathbf{E}\left[\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{x}_{t-1},\boldsymbol{c}_{t})|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t}\right]|\Lambda_{T}\right]$ $\leq\frac{\beta}{2}\sum_{t=1}^{T}\mathbf{E}\left[\left\lVert P_{t}\boldsymbol{\zeta}_{t}\right\rVert^{2}\right]$
$\leq\frac{\beta}{2}\sum_{t=1}^{T}\mathbf{E}\left[\left\lVert P_{t}\right\rVert^{2}_{\text{op}}\right]\leq\frac{\beta}{2\alpha}\sum_{t=1}^{T}\frac{1}{\eta_{t}t},$
where $\left\lVert\cdot\right\rVert_{\text{op}}$ denotes the operator norm. By the fact that $f(\cdot,\boldsymbol{c}_{t})$ is a convex function and by Jensen’s inequality, we deduce that term II is negative. For term III, by Assumption 2(ii) we have
term III $\leq\frac{\beta}{2}\sum_{i=1}^{T}\mathbf{E}\left[\left\lVert P_{t}\boldsymbol{U}\right\rVert^{2}\right]\leq\frac{\beta}{2}\sum_{t=1}^{T}\mathbf{E}\left[\left\lVert P_{t}\right\rVert^{2}_{\text{op}}\right]\leq\frac{\beta}{2\alpha}\sum_{t=1}^{T}\frac{1}{\eta_{t}t}.$
To control term IV, first note that by Lemma 16 $\{{\sf f}_{t,P}(\cdot,\boldsymbol{c}_{t})\}_{t=1}^{T}$ are $\alpha$-strongly convex functions. Furthermore, by Lemma 17 we have that $\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t-1},\boldsymbol{c}_{t},\Lambda_{T}\right]=\nabla{\sf f}_{t,P}(\boldsymbol{x}_{t-1},\boldsymbol{c}_{t})$. Therefore, by Lemma 15 we obtain
term IV $\leq 2d^{2}q_{T}^{2}\sum_{t=1}^{T}\eta_{t}+4dq_{T}\sqrt{\mu T}+\eta_{T}^{-1}\mathcal{R}(\boldsymbol{z}^{\prime}),$
where $q_{T}=M+2\sigma\sqrt{\log(T+1)}$ and $\boldsymbol{z}^{\prime}=(1-1/T)\boldsymbol{z}+(1/T)\boldsymbol{x}_{0}$. By Lemma 18 we have $\mathcal{R}(\boldsymbol{z}^{\prime})\leq\mu\log(T)$, so that
term IV $\leq 2d^{2}q_{T}^{2}\sum_{t=1}^{T}\eta_{t}+4dq_{T}\sqrt{\mu T}+\eta_{T}^{-1}\mu\log(T).$
Thus,
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]\leq 2d^{2}q_{T}^{2}\sum_{t=1}^{T}\eta_{t}+4dq_{T}\sqrt{\mu T}+\eta_{T}^{-1}\mu\log(T)+\frac{\beta}{\alpha}\sum_{t=1}^{T}\frac{1}{\eta_{t}t}.$
Since $\eta_{t}=(4dq_{T})^{-1}\min\big(1,\frac{\nu\log(t+1)}{t}\big)^{\frac{1}{2}}$ with $\nu=16(\mu+\beta/\alpha)$ we get
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]$ $\leq dq_{T}\sqrt{\nu\log(T+1)T}+4dq_{T}\sqrt{\mu T}+$
$+4dq_{T}\mu\max\left(\log^{2}(T+1),\frac{T\log(T+1)}{\nu}\right)^{\frac{1}{2}}+\frac{\beta}{\alpha}\sum_{t=1}^{T}\frac{1}{\eta_{t}t}.$
Using the fact that $16\mu\leq\nu$, and $\max\{a,b\}\leq a+b$ for $a,b>0$, we further simplify the above bound to obtain
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]$ $\leq\frac{9}{4}dq_{T}\sqrt{\nu\log(T+1)T}+4dq_{T}\mu\log(T+1)+\frac{\beta}{\alpha}\sum_{t=1}^{T}\frac{1}{\eta_{t}t}.$
Since $t/\log(t+1)\geq\sqrt{t}$ for all $t\in[T]$, then if $t\geq\nu^{2}$ we have $t\geq\nu\log(t+1)$. Thus,
$\frac{\beta}{\alpha}\sum_{t=1}^{T}\frac{1}{\eta_{t}t}$ $\leq\frac{4\beta}{\alpha}dq_{T}\left(\sum_{t=1}^{\nu^{2}}\frac{1}{t}+\sum_{t=2}^{T}\sqrt{\frac{1}{\nu t\log(t+1)}}\right)$
$\leq 4dq_{T}\left(\frac{\beta}{\alpha}\log(\nu^{2}+1)+\frac{2\beta}{\alpha}\sqrt{\frac{T}{\nu}}\right)\leq 4dq_{T}\left(\frac{\beta}{\alpha}\log(\nu^{2}+1)+\frac{1}{8}\sqrt{\nu T}\right),$
where the last inequality is due to the fact that $16\beta/\alpha\leq\nu$. Thus,
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z}^{\prime},\boldsymbol{c}_{t})|\Lambda_{T}\right]$ $\leq dq_{T}\left(\frac{11}{4}\sqrt{\nu\log(T+1)T}+4\left(\mu\log(T+1)+\frac{\beta}{\alpha}\log(\nu^{2}+1)\right)\right)$
$\leq\frac{dq_{T}}{4}\left(11\sqrt{\nu\log(T+1)T}+\nu\left(\log(T+1)+\log(\nu^{2}+1)\right)\right).$
Since $q_{T}=M+2\sigma\sqrt{\log(T+1)}$ we have
[…]
$\quad+\nu\left(\log(T+1)+2\log(\nu+1)\right)\bigg).$
Combining (38), (39), and (40) implies
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z},\boldsymbol{c}_{t})|\Lambda_{T}\right]$ $\leq\frac{d}{4}\big(M+2\sigma\sqrt{\log(T+1)}\big)\bigg(11\sqrt{\nu T\log(T+1)}$
$\quad+\nu\left(\log(T+1)+2\log(\nu+1)\right)\bigg)+6M.$
Note that since $\nu\geq 16$ and $2\log(\nu+1)\geq 5.5$, we have
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z},\boldsymbol{c}_{t})|\Lambda_{T}\right]$ $\leq\frac{d}{4}\big(5.3M+2\sigma\sqrt{\log(T+1)}\big)\bigg(11\sqrt{\nu T\log(T+1)}$
$\quad+\nu\left(\log(T+1)+2\log(\nu+1)\right)\bigg).$
Thus, we can write
$\sum_{t=1}^{T}\mathbf{E}\left[f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-f(\boldsymbol{z},\boldsymbol{c}_{t})\right]$ $\leq F(T)F_{1}(T),$
where
$F(T)=11\sqrt{\nu T\log(T+1)}+\nu\left(\log(T+1)+2\log(\nu+1)\right),$
and
$F_{1}(T)=\frac{d}{4}\left(5.3M+2\sigma\sqrt{\log(T+1)}\right).$
Noticing that $F$ is a concave and $F_{1}$ is a non-decreasing function we complete the first part of the proof. For the second part of the proof, assume that $T\geq 16(\mu+\beta/\alpha)^{2}$. In this case we have $T\geq\nu$ since $\beta/\alpha\geq 1$. Therefore,
$F(T)\leq 11\sqrt{\nu T\log(T+1)}+3\nu\log(T+1).$
On the other hand, the condition $T\geq 16(\mu+\beta/\alpha)^{2}$ implies that
$\sqrt{\nu T\log(T+1)}\geq\frac{\nu}{2}T^{\frac{1}{4}}\sqrt{\log(T+1)}\geq\frac{\nu}{2}\log(T+1),$
and therefore,
$F(T)$ $\leq 17\sqrt{\nu T\log(T+1)}=68\sqrt{\left(\mu+\frac{\beta}{\alpha}\right)T\log(T+1)}.$
Proof [Proof of Corollary 9] Since $F$ in (13) is a concave function we can apply (10) to obtain
$\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\big)\big]$ $\leq K^{p}F\left(\frac{T}{K^{p}}\right)F_{1}(T)+2LT\left(\frac{\sqrt{p}}{K}\right)^{\gamma}.$
Using the definition of $F$ and the inequality $\mu\leq\texttt{A}d^{\rho}$, we get
$\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\big)\big]$ $\leq C_{*}\Big(L_{1}d^{1+\rho}K^{p}+L_{2}\sqrt{K^{p}d^{2+\rho}T}+LT\left(\frac{\sqrt{p}}{K}\right)^{\gamma}\Big),$ (41)
where $L_{1}=\sqrt{\log(T+1)}\log(Td+1)$, $L_{2}=\log(T+1)$, and $C_{*}>0$ is a constant depending only on $M,\sigma,\beta/\alpha$.
Consider now two cases. First, let $Lp^{\frac{\gamma}{2}}d^{-(1+\rho/2)}T^{1/2}\log^{-1}(T+1)>1$. In this case, since $x/2\leq\lfloor x\rfloor\leq x$ for $x>1$, we obtain that
$K=K(1+\rho/2,1/2,1,\gamma)\asymp\Big(Lp^{\frac{\gamma}{2}}d^{-(1+\rho/2)}T^{1/2}\log^{-1}(T+1)\Big)^{\frac{2}{p+2\gamma}},$
so that, to within a constant depending only on $M,\sigma,\beta/\alpha$, the right hand side of (41) is of the order
$L_{1}d^{1+\rho}\Big(Lp^{\frac{\gamma}{2}}d^{-(1+\rho/2)}L_{2}^{-1}\Big)^{\frac{2p}{p+2\gamma}}T^{\frac{p}{p+2\gamma}}+L_{2}^{\frac{2\gamma}{p+2\gamma}}\Big(Lp^{\frac{\gamma}{2}}\Big)^{\frac{p}{p+2\gamma}}d^{\frac{(2+\rho)\gamma}{p+2\gamma}}T^{\frac{p+\gamma}{p+2\gamma}}.$
In the second case, $Lp^{\frac{\gamma}{2}}d^{-(1+\rho/2)}T^{1/2}\log^{-1}(T+1)\leq 1$. Then $K(1+\rho/2,1/2,1,\gamma)=1$, so that the right hand side of (41) does not exceed
$C_{*}(L_{1}d^{1+\rho}+L_{2}\sqrt{d^{2+\rho}T}+(Lp^{\frac{\gamma}{2}}T^{1/2})\sqrt{T})\leq C_{*}(L_{1}d^{1+\rho}+(L_{2}+1)\sqrt{d^{2+\rho}T}).$
Putting these remarks together we get
$\mathbf{E}\big[\sum_{t=1}^{T}\big(f(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{z}\in\Theta}f(\boldsymbol{z},\boldsymbol{c}_{t})\big)\big]\leq C\max\bigg\{L_{1}d^{1+\rho}+L_{2}\sqrt{d^{2+\rho}T},\,$
$\qquad\qquad\qquad\qquad\qquad L_{1}\Big(Lp^{\frac{\gamma}{2}}L_{2}^{-1}\Big)^{\frac{2p}{p+2\gamma}}d^{\frac{2\gamma(1+\rho)-p}{p+2\gamma}}T^{\frac{p}{p+2\gamma}},\,L_{2}^{\frac{2\gamma}{p+2\gamma}}\Big(Lp^{\frac{\gamma}{2}}\Big)^{\frac{p}{p+2\gamma}}d^{\frac{(2+\rho)\gamma}{p+2\gamma}}T^{\frac{p+\gamma}{p+2\gamma}}\bigg\},$
where $C>0$ is a constant depending only on $M,\sigma,\beta/\alpha$.
Appendix C Proof of Theorem 11
Lemma 19
Suppose that $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are two probability measures that are defined on a probability space $(X,\mathcal{M})$, such that $\mathbf{P}_{1}$ is absolutely continuous with respect to $\mathbf{P}_{2}$. Let $\Upsilon\in\mathcal{M}$ be such that $\mathbf{P}_{1}(\Upsilon)=\mathbf{P}_{2}(\Upsilon)\neq 0$. Let $\tilde{\mathbf{P}}_{1}$ and $\tilde{\mathbf{P}}_{2}$ be the conditional probability measures of $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ given the event $\Upsilon$, respectively, i.e., for $i=1,2$ and $\Gamma\in\mathcal{M}$
$\tilde{\mathbf{P}}_{i}(\Gamma)=\mathbf{P}_{i}(\Gamma\cap\Upsilon)\mathbf{P}_{i}^{-1}(\Upsilon).$
Then,
  1. 1.
    For $i=1,2$
    $\frac{\,\mathrm{d}\tilde{\mathbf{P}}_{i}(\boldsymbol{x})}{\,\mathrm{d}\mathbf{P}_{i}(\boldsymbol{x})}=\mathbb{1}(\boldsymbol{x}\in\Upsilon)\mathbf{P}_{i}^{-1}(\Upsilon).$
  2. 2.
    $KL(\tilde{\mathbf{P}}_{1},\tilde{\mathbf{P}}_{2})=\int\log\left(\frac{\,\mathrm{d}\mathbf{P}_{1}(\boldsymbol{x})}{\,\mathrm{d}\mathbf{P}_{2}(\boldsymbol{x})}\right)\mathbb{1}(\boldsymbol{x}\in\Upsilon)\mathbf{P}_{1}^{-1}(\Upsilon)\,\mathrm{d}\mathbf{P}_{1}(\boldsymbol{x}).$
Proof [Proof of Theorem 11] We choose a finite set of functions in a way that its members cannot be distinguished from each other with positive probability but are separated enough from each other to guarantee that the maximal regret for this family is not smaller than the desired lower bound.
Set $K=\max\left(1,\left\lfloor\left(\min\left(1,L^{2}\right)T\right)^{\frac{1}{p+2\gamma}}\right\rfloor\right)$. Let $(B_{i})_{i=1}^{K^{p}}$ be a partition of $0,1]^{p}$, where $B_{i}$’s are disjoint cubes with edges of length $1/K$. Then, for any $\boldsymbol{c}\in[0,1]^{p}$ there exists $i\in\{1,\dots,K^{p}\}$ that satisfies $d(\boldsymbol{c},\partial B_{i})\leq\delta$, where $\partial B_{i}$ is the boundary of $B_{i}$, $d(\boldsymbol{c},\partial B_{i})=\min_{\boldsymbol{u}\in\partial B_{i}}\left\lVert\boldsymbol{c}-\boldsymbol{u}\right\rVert$ and $\delta=\frac{1}{2K}$, cf. (46).
Let $\eta_{0}:\mathbb{R}\to\mathbb{R}$ be an infinitely many times differentiable function such that
$\eta_{0}(x)=\begin{cases}=1\quad\quad&\text{if }|x|\leq 1/4,\\ \in(0,1)\quad\quad&\text{if }1/4<|x|<1,\\ =0\quad\quad&\text{if }|x|\geq 1.\end{cases}$
Set $\eta(x)=\int_{-\infty}^{x}\eta_{0}(u)\,\mathrm{d}u$. Let $\Omega=\{-1,1\}^{K^{p}}$ and $\{-1,1\}^{d-1}$ be the sets of binary sequences of length $K^{p}$ and d1d-1, respectively. Consider the finite set of functions $f_{\boldsymbol{\omega},\boldsymbol{\tau}}:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$, $\boldsymbol{\omega}\in\Omega$, $\boldsymbol{\tau}\in\{-1,1\}^{d-1}$ such that:
$f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})=\alpha\left\lVert\boldsymbol{x}\right\rVert^{2}+r_{1}L\eta\Big(x_{1}\delta^{-\frac{\gamma}{2}}\Big)$ $\left(\sum_{j=1}^{K^{p}}\omega_{j}d^{\gamma}(\boldsymbol{c},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}\in B_{j}\right)\right)$ (42)
$+r_{2}h^{2}\left(\sum_{i=2}^{d}\tau_{i}\eta\left(\frac{x_{i}}{h}\right)\right),\quad\quad\boldsymbol{x}=(x_{1},\dots,x_{d}),$
where $\omega_{j},\tau_{i}\in\{-1,1\}$, $h=\min\left(d^{-\frac{1}{2}},T^{-\frac{1}{4}}\right)$ and $r_{1},r_{2}$ are fixed positive numbers that will be chosen small enough. By Lemma 21 we have that $f_{\boldsymbol{\omega},\boldsymbol{\tau}}\in\mathcal{F}_{\alpha,\beta}(M)\cap\mathcal{F}_{\gamma}(L)$ provided $r_{1}\leq\min\left(1,1/L^{2}\right)\min\left(1/2\eta(1),\alpha/L^{\prime}\right),r_{2}\leq\min\left(1/2\eta(1),\alpha/L^{\prime}\right)$, where $L^{\prime}=\max_{x\in\mathbb{R}}|\eta^{\prime\prime}\left(x\right)|$. Furthermore, by choosing $r_{1}\leq\alpha/\left(2\max\left(1,L\right)\right)$ and $r_{2}\leq\alpha/2$ we have
$\alpha^{-1}\delta^{-\gamma}r_{1}L\sum_{j=1}^{K^{p}}\omega_{j}d^{\gamma}(\boldsymbol{c}_{t},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}\in B_{j}\right)\leq\frac{1}{2},\quad\quad\text{}\quad\quad\alpha^{-1}r_{2}\leq\frac{1}{2},$
and that under this condition the equation $\nabla_{\boldsymbol{x}}f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})=0$ has the solution
$\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c})=\left(x^{*}_{1}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}),\dots,x^{*}_{d}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c})\right),$
where $x^{*}_{1}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c})=-(1/2)\alpha^{-1}\delta^{-\frac{\gamma}{2}}r_{1}L\sum_{j=1}^{K^{p}}\omega_{j}d^{\gamma}(\boldsymbol{c}_{t},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}\in B_{j}\right)$ and $x^{*}_{i}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c})=-(1/2)\tau_{i}r_{2}\alpha^{-1}h$, for $2\leq i\leq d$. Furthermore,
$\left\lVert\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c})\right\rVert^{2}\leq\frac{1}{4\alpha^{2}}\left(r_{1}^{2}L^{2}\delta^{\gamma}+dr_{2}^{2}h^{2}\right)\leq\frac{1}{8},$
which implies $\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c})\in\Theta$.
For any fixed $\boldsymbol{\omega}\in\Omega,\boldsymbol{\tau}\in\{-1,1\}^{d-1}$, and $1\leq t\leq T$, we denote by $\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$ the probability measure corresponding to the joint distribution of $(\boldsymbol{z}_{1},y_{1}^{t},\boldsymbol{c}_{1}^{t})$ where $y_{k}=f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{z}_{k},\boldsymbol{c}_{k})+\xi_{k}$ with independent identically distributed $\xi_{k}$’s such that (15) holds and $\boldsymbol{z}_{k}$’s chosen by the randomized procedure $\pi$. We have
$\,\mathrm{d}\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left(\boldsymbol{z}_{1},y_{1}^{t},\boldsymbol{c}_{1}^{t}\right)=\,\mathrm{d}F\left(y_{1}-f_{\boldsymbol{\omega},\boldsymbol{\tau}}\left(\boldsymbol{z}_{1},\boldsymbol{c}_{1}\right)\right)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{1})\prod_{i=2}^{t}\,\mathrm{d}F\left(y_{i}-f_{\boldsymbol{\omega},\boldsymbol{\tau}}\left(\Phi_{i}\left(\boldsymbol{z}_{1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i}\right),\boldsymbol{c}_{i}\right)\right)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i}).$ (43)
Without loss of generality, we omit here the dependence of $\Phi_{i}$ on $\boldsymbol{z}_{2},\dots,\boldsymbol{z}_{i-1}$ since $\boldsymbol{z}_{i},i\geq 2$, is a Borel function of $\boldsymbol{z}_{1},y_{1},\dots,y_{i-1},\boldsymbol{c}_{1},\dots,\boldsymbol{c}_{i}$. Let $\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$ denote the expectation w.r.t. $\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$.
Note that
$\sum_{t=1}^{T}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{z}_{t},\boldsymbol{c}_{t})-\min_{\boldsymbol{x}\in\Theta}f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}_{t})\right]\geq\frac{\alpha}{2}\sum_{t=1}^{T}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert^{2}\right]$ (44)
due to the uniform strong convexity of functions $\boldsymbol{x}\mapsto f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})$.
Consider the statistics
$(\hat{\boldsymbol{\omega}}_{t},\hat{\boldsymbol{\tau}}_{t})\in\operatorname*{arg\,min}_{\boldsymbol{\omega}\in\Omega,\boldsymbol{\tau}\in\{-1,1\}^{d-1}}\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert.$
Since $\left\lVert\boldsymbol{x}^{*}(\hat{\boldsymbol{\omega}}_{t},\hat{\boldsymbol{\tau}}_{t},\boldsymbol{c}_{t})-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert\leq\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\hat{\boldsymbol{\omega}}_{t},\hat{\boldsymbol{\tau}}_{t},\boldsymbol{c}_{t})\right\rVert+\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert\leq 2\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert$ for all $\boldsymbol{\omega}\in\Omega,\boldsymbol{\tau}\in\{-1,1\}^{d-1}$ we obtain
$\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert^{2}\right]$ $\geq\frac{1}{4}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\left\lVert\boldsymbol{x}^{*}(\hat{\boldsymbol{\omega}}_{t},\hat{\boldsymbol{\tau}}_{t},\boldsymbol{c}_{t})-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert^{2}\right]$
$=\frac{1}{4}\alpha^{-2}L^{2}\delta^{-\gamma}r_{1}^{2}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\sum_{j=1}^{K^{p}}d^{2\gamma}(\boldsymbol{c}_{t},\partial B_{j})\mathbb{1}\left(\hat{\omega}_{t,j}\neq\omega_{j}\right)\mathbb{1}\left(\boldsymbol{c}_{t}\in B_{j}\right)\right]$
[…]
where $\rho(\boldsymbol{\tau},\hat{\boldsymbol{\tau}}_{t})=\sum_{i=2}^{d}\mathbb{1}\left(\boldsymbol{\tau}_{i}\neq\hat{\boldsymbol{\tau}}_{t,i}\right)$ is the Hamming distance between $\boldsymbol{\tau}$ and $\hat{\boldsymbol{\tau}}_{t}$, with $\hat{\boldsymbol{\omega}}_{t}=\left(\hat{\omega}_{t,1},\dots,\hat{\omega}_{t,K^{p}}\right)$ and $\hat{\boldsymbol{\tau}}_{t}=\left(\hat{\tau}_{t,2},\dots,\hat{\tau}_{t,d}\right)$. Note that, due to (47), for $\varepsilon=1/2$ we have
$\mathbb{P}_{K}\left[d(\boldsymbol{c}_{t},\partial B_{j})\geq\frac{1}{4K}\Big|\boldsymbol{c}_{t}\in B_{j}\right]\geq\mathbb{P}_{K}\left[\left\lVert\boldsymbol{c}_{t}-\boldsymbol{b}_{j}\right\rVert_{\infty}\leq\frac{1}{4K}\Big|\boldsymbol{c}_{t}\in B_{j}\right]=1,$
which yields
$\mathbb{P}_{K}\left[d^{2\gamma}(\boldsymbol{c}_{t},\partial B_{j})\geq\left(\frac{\delta^{2}}{4p^{2}}\right)^{\gamma}\Big|\boldsymbol{c}_{t}\in B_{j}\right]=1.$
Thus,
$\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$ $\left[\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{c}_{t})\right\rVert^{2}\right]\geq\frac{1}{4}\alpha^{-2}\delta^{-\gamma}r_{1}^{2}L^{2}\sum_{j=1}^{K^{p}}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[d^{2\gamma}(\boldsymbol{c}_{t},\partial B_{j})\mathbb{1}\left(\hat{\omega}_{t,j}\neq\omega_{j}\right)|\boldsymbol{c}_{t}\in B_{j}\right]\mathbb{P}_{K}\left(\boldsymbol{c}_{t}\in B_{j}\right)$
[…]
$=\frac{1}{4}\alpha^{-2}r_{1}^{2}L^{2}\delta^{-\gamma}\sum_{j=1}^{K^{p}}\bigg(\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[d^{2\gamma}(\boldsymbol{c}_{t},\partial B_{j})\mathbb{1}\left(\hat{\omega}_{t,j}\neq\omega_{j}\right)\Big|\boldsymbol{c}_{t}\in B_{j},d^{2\gamma}(\boldsymbol{c}_{t},\partial B_{j})\geq\left(\frac{\delta^{2}}{4p^{2}}\right)^{\gamma}\right]$
[…]
$\geq\frac{\alpha^{-2}}{4^{\gamma+1}}r_{1}^{2}L^{2}\delta^{\gamma}\sum_{j=1}^{K^{p}}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq\omega_{j}\right)\Big|\boldsymbol{c}_{t}\in B_{j},d^{2\gamma}(\boldsymbol{c}_{t},\partial B_{j})\geq\left(\frac{\delta^{2}}{4p^{2}}\right)^{\gamma}\right]\mathbb{P}_{K}\left(\boldsymbol{c}_{t}\in B_{j}\right)$
[…]
Introduce the random event $\Upsilon_{t,j}=\{\boldsymbol{c}_{t}\in B_{j},\ d(\boldsymbol{c}_{t},\partial B_{j})\geq\frac{\delta}{2p}\}$. Summing both sides from $1$ to $T$, taking the maximum over $\Omega$ and $\{-1,1\}^{d-1}$ and then taking the minimum over all statistics $\hat{\boldsymbol{\omega}}_{t}$ with values in $\Omega$ and $\hat{\boldsymbol{\tau}}_{t}$ with values in $\{-1,1\}^{d-1}$ we obtain
$\min_{\hat{\boldsymbol{\omega}}_{1},\dots,\hat{\boldsymbol{\omega}}_{T}}$ $\max_{\boldsymbol{\omega}\in\Omega}\sum_{t=1}^{T}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{c}_{t})\right\rVert^{2}\right]\geq\frac{\alpha^{-2}}{4^{\gamma+1}}r_{1}^{2}L^{2}\delta^{\gamma}K^{-p}\underbrace{\min_{\hat{\boldsymbol{\omega}}_{1},\dots,\hat{\boldsymbol{\omega}}_{T}}\max_{\boldsymbol{\omega}\in\Omega}\sum_{t=1}^{T}\sum_{j=1}^{K^{p}}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq\omega_{j}\right)|\Upsilon_{t,j}\right]}_{\text{term I}}$
$\phantom{000000}+\frac{1}{4}\alpha^{-2}r_{2}^{2}h^{2}\underbrace{\min_{\hat{\boldsymbol{\tau}}_{1},\dots,\hat{\boldsymbol{\tau}}_{T}}\max_{\boldsymbol{\omega}\in\Omega,\boldsymbol{\tau}\in\{-1,1\}^{d-1}}\sum_{t=1}^{T}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\rho(\boldsymbol{\tau},\hat{\boldsymbol{\tau}_{t}})\right]}_{\text{term II}}.$
We will treat terms I and II separately. For term I we can write
term I $\geq 2^{-K^{p}-d+1}\min_{\hat{\boldsymbol{\omega}}_{1},\dots,\hat{\boldsymbol{\omega}}_{T}}\sum_{t=1}^{T}\sum_{j=1}^{K^{p}}\sum_{\boldsymbol{\tau}\in\{-1,1\}^{d-1}}\sum_{\boldsymbol{\omega}\in\Omega}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq\omega_{j}\right)|\Upsilon_{t,j}\right]$
$\geq 2^{-K^{p}-d+1}\sum_{t=1}^{T}\sum_{j=1}^{K^{p}}\sum_{\boldsymbol{\tau}\in\{-1,1\}^{d-1}}\sum_{\boldsymbol{\omega}\in\Omega}\min_{\hat{\boldsymbol{\omega}}_{t,j}}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq\omega_{j}\right)|\Upsilon_{t,j}\right]$
Let $\Omega_{j}=\{\boldsymbol{\omega}\in\Omega:\omega_{j}=1\}$, and for any $\boldsymbol{\omega}\in\Omega_{j}$, denote $\bar{\boldsymbol{\omega}}$ such that $\omega_{i}=\bar{\omega}_{i}$, for any $i\neq j$, and $\bar{\omega}_{j}=-1$. Thus,
term I $\geq 2^{-K^{p}-d+1}\sum_{t=1}^{T}\sum_{j=1}^{K^{p}}\sum_{\boldsymbol{\tau}\in\{-1,1\}^{d-1}}\sum_{\boldsymbol{\omega}\in\Omega_{j}}\min_{\hat{\omega}_{t,j}}\left(\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq 1\right)|\Upsilon_{t,j}\right]+\mathbf{E}_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq-1\right)|\Upsilon_{t,j}\right]\right)$
$\geq\frac{1}{2}\sum_{t=1}^{T}\sum_{j=1}^{K^{p}}\min_{\boldsymbol{\tau}\in\{-1,1\}^{d-1},\boldsymbol{\omega}\in\Omega_{j}}\min_{\hat{\omega}_{t,j}}\left(\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq 1\right)|\Upsilon_{t,j}\right]+\mathbf{E}_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau},t}\left[\mathbb{1}\left(\hat{\omega}_{t,j}\neq-1\right)|\Upsilon_{t,j}\right]\right).$
For any $\boldsymbol{\omega}\in\Omega_{j}$ and $\boldsymbol{\tau}\in\{-1,1\}^{d-1}$, let $\tilde{\mathbf{P}}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$ and $\tilde{\mathbf{P}}_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau},t}$ be the conditional probability distributions of $\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$ and $\mathbf{P}_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau},t}$ on the event $\Upsilon_{t,j}$, respectively. Then, by Lemma 20 we have
$KL\bigg(\tilde{\mathbf{P}}_{\boldsymbol{\omega},\boldsymbol{\tau},t},\tilde{\mathbf{P}}_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau},t}\bigg)$ $\leq I_{0}\sum_{i=1}^{t-1}\int\max_{\boldsymbol{x}\in\Theta}|f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}_{i})-f_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}_{i})|^{2}\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i})$
$\phantom{0000}+I_{0}\int\max_{\boldsymbol{x}\in\Theta}|f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}_{t})-f_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}_{t})|^{2}\mathbb{1}(\boldsymbol{c}_{t}\in\Upsilon_{t,j})\mathbb{P}_{K}^{-1}(\Upsilon_{t,j})\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{t})$
$\leq I_{0}\sum_{i=1}^{t-1}\int|2r_{1}L\eta(1)d^{\gamma}(\boldsymbol{c}_{i},\partial B_{j})|^{2}\mathbb{1}\left(\boldsymbol{c}_{i}\in B_{j}\right)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i})$
$\phantom{0000}+I_{0}\int|2r_{1}\eta(1)d^{\gamma}(\boldsymbol{c}_{t},\partial B_{j})|^{2}\mathbb{1}(\boldsymbol{c}_{t}\in\Upsilon_{t,j})\mathbb{P}_{K}^{-1}(\Upsilon_{t,j})\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{t})$
$\leq 4I_{0}r_{1}^{2}L^{2}\eta^{2}(1)\delta^{2\gamma}\left(\sum_{i=1}^{t-1}\int\mathbb{1}\left(\boldsymbol{c}_{i}\in B_{j}\right)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i})+1\right)\leq 4I_{0}r_{1}^{2}L^{2}\eta^{2}(1)\delta^{2\gamma}\left(tK^{-p}+1\right).$
Since $\delta=1/(2K)$, $K=\max\left(1,\left\lfloor\left(\min\left(1,L^{2}\right)T\right)^{\frac{1}{p+2\gamma}}\right\rfloor\right)$ we obtain that
$KL\left(\tilde{\mathbf{P}}_{\boldsymbol{\omega},\boldsymbol{\tau},t},\tilde{\mathbf{P}}_{\bar{\boldsymbol{\omega}},\boldsymbol{\tau},t}\right)\leq\log(2)$
if we choose $r_{1}\leq\left(\log(2)/(8I_{0}\eta^{2}(1)\max(1,L^{2}))\right)^{1/2}$. Therefore, using Theorem 2.12 from Tsybakov (2009) we get
$\text{term I}\geq\frac{TK^{p}}{4}\exp(-\log(2))=\frac{TK^{p}}{8}.$
Moreover, for any $\boldsymbol{\omega}\in\Omega$ and $\boldsymbol{\tau},\boldsymbol{\tau}^{\prime}\in\{-1,1\}^{d-1}$ such that $\rho(\boldsymbol{\tau},\boldsymbol{\tau}^{\prime})=1$ we have
$KL\left(\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau},t},\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau}^{\prime},t}\right)$ $=\int\log\left(\frac{\,\mathrm{d}\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau},t}}{\,\mathrm{d}\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau}^{\prime},t}}\right)\,\mathrm{d}\mathbf{P}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$
$=\int\bigg[\log\left(\frac{\,\mathrm{d}F(y_{1}-f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{z}_{1},\boldsymbol{c}_{1}))}{\,\mathrm{d}F(y_{1}-f_{\boldsymbol{\omega},\boldsymbol{\tau}^{\prime}}(\boldsymbol{z}_{1},\boldsymbol{c}_{1}))}\right)+$
$\phantom{0000000}+\sum_{i=2}^{t}\log\left(\frac{\,\mathrm{d}F(y_{i}-f_{\boldsymbol{\omega},\boldsymbol{\tau}}\left(\Phi_{i}\left(\boldsymbol{z}_{1}^{i-1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i}\right),\boldsymbol{c}_{i})\right)}{\,\mathrm{d}F(y_{i}-f_{\boldsymbol{\omega},\boldsymbol{\tau}^{\prime}}\left(\Phi_{i}\left(\boldsymbol{z}_{1}^{i-1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i}\right),\boldsymbol{c}_{i})\right)}\right)\bigg]$
$\phantom{000000000}\,\mathrm{d}F\left(y_{i}-f_{\boldsymbol{\omega},\boldsymbol{\tau}}\left(\boldsymbol{z}_{1}^{i-1},\boldsymbol{c}_{1}\right)\right)\prod_{i=2}^{t}\,\mathrm{d}F\left(y_{i}-f_{\boldsymbol{\omega},\boldsymbol{\tau}}\left(\Phi_{i}\left(\boldsymbol{z}_{1}^{i-1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i}\right),\boldsymbol{c}_{i}\right)\right)$
$\leq I_{0}\sum_{i=1}^{t}\max_{\boldsymbol{x}\in\Theta}|f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}_{i})-f_{\boldsymbol{\omega},\boldsymbol{\tau}^{\prime}}(\boldsymbol{x},\boldsymbol{c}_{i})|^{2}=4tI_{0}r_{2}^{2}h^{4}\eta^{2}(1).$
Using the fact that $h\leq T^{-\frac{1}{4}}$, choosing $r_{2}\leq\left(\log(2)/(4I_{0}\eta^{2}(1))\right)^{1/2}$ and applying again Theorem 2.12 from Tsybakov (2009) implies
$\text{term II}\geq\frac{T(d-1)}{4}\exp(-\log(2))\geq\frac{Td}{16}.$
Putting together the bounds for terms I and II we get
$\min_{\hat{\boldsymbol{\omega}}_{1},\hat{\boldsymbol{\tau}}_{1}\dots,\hat{\boldsymbol{\omega}}_{T},\hat{\boldsymbol{\tau}}_{T}}\max_{\boldsymbol{\omega}\in\Omega,\tau\in\{-1,1\}^{d-1}}\sum_{t=1}^{T}\mathbf{E}_{\boldsymbol{\omega},\boldsymbol{\tau},t}$ $\left[\left\lVert\boldsymbol{z}_{t}-\boldsymbol{x}^{*}(\boldsymbol{\omega},\boldsymbol{\tau},\boldsymbol{c}_{t})\right\rVert^{2}\right]\geq\frac{\alpha^{-2}}{256}r_{1}^{2}T\delta^{\gamma}+\frac{\alpha^{-2}}{64}r_{2}^{2}dTh^{2}$
$\geq\frac{\alpha^{-2}}{2048}r_{3}^{2}\min\left(1,L^{\frac{2(d+\gamma)}{d+2\gamma}}\right)T^{\frac{d+\gamma}{d+2\gamma}}+\frac{\alpha^{-2}}{64}r_{2}^{2}\min\left(T,dT^{\frac{1}{2}}\right),$ (45)
where $r_{3}=r_{2}/\min(1,L)$. Considering the fact that $r_{2}$ and $r_{3}$ are independent of $d,p,L$, and $T$, (45) combined with (44) implies the theorem.
Recall the definition of $\mathbb{P}_{K}$ in (16). The following figure illustrates the support of $\mathbb{P}_{K}$ for $K=3$ and $p=2$. By definition, $\mathbb{P}_{K}$ yields a uniform distribution for each black square in the figure.
Figure 1: Support of $\mathbb{P}_{K}$ for $K=3$ and $p=2$
The next lemma provides an upper bound on the Kullback-Leibler divergence of two probability measures, which plays a key role in our analysis of the lower bound.
Lemma 20
With the same notation as in the proof of Theorem 11, for $\boldsymbol{\omega},\boldsymbol{\omega}^{\prime}\in\Omega$ and $t\geq 1$ denote $\mathbf{P}_{\boldsymbol{\omega},t}$ and $\mathbf{P}_{\boldsymbol{\omega}^{\prime},t}$ as they are defined in (43). Let $\Upsilon\subseteq[0,1]^{p}$ such that $\mathbb{P}_{K}(\Upsilon)\neq 0$. Let $\tilde{\mathbf{P}}_{\boldsymbol{\omega},t}$ $\tilde{\mathbf{P}}_{\boldsymbol{\omega}^{\prime},t}$ be the conditional probability measures of ${\mathbf{P}}_{\boldsymbol{\omega},t}$ ${\mathbf{P}}_{\boldsymbol{\omega}^{\prime},t}$ given the event $\boldsymbol{c}_{t}\in\Upsilon$, respectively. Then, if (15) holds, we have
$KL(\tilde{\mathbf{P}}_{\boldsymbol{\omega},t},\tilde{\mathbf{P}}_{\boldsymbol{\omega}^{\prime},t})$ $\leq I_{0}\sum_{i=1}^{t-1}\max_{\boldsymbol{x}\in\Theta}|f_{\boldsymbol{\omega},t}(\boldsymbol{x},\boldsymbol{c}_{i})-f_{\boldsymbol{\omega}^{\prime},t}(\boldsymbol{x},\boldsymbol{c}_{i})|\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i})$
$\phantom{0000}+I_{o}\int\max_{\boldsymbol{x}\in\Theta}|f_{\boldsymbol{\omega},t}(\boldsymbol{x},\boldsymbol{c}_{i})-f_{\boldsymbol{\omega}^{\prime},t}(\boldsymbol{x},\boldsymbol{c}_{i})|\mathbb{1}(\boldsymbol{c}_{t}\in\Upsilon)\mathbb{P}_{K}^{-1}(\Upsilon)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{t}).$
Proof By Lemma 19(i), for any $\boldsymbol{\omega}\in\Omega$ we have
$\,\mathrm{d}\tilde{\mathbf{P}}_{\boldsymbol{\omega},t}(\boldsymbol{z}_{1},y_{1}^{t},\boldsymbol{c}_{1}^{t})$ $=\,\mathrm{d}F(y_{1}-f_{\boldsymbol{\omega}}(\boldsymbol{z}_{1},\boldsymbol{c}_{1}))\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{1})\left(\prod_{i=2}^{t-1}\,\mathrm{d}F(y_{i}-f_{\boldsymbol{\omega}}(\pi_{i}(\boldsymbol{z}_{1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i-1})))\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i})\right)$
$\phantom{00000}\,\mathrm{d}F(y_{t}-f_{\boldsymbol{\omega}}(\pi_{i}(\boldsymbol{z}_{1},y_{1}^{t-1},\boldsymbol{c}_{1}^{t})))\mathbb{1}(\boldsymbol{c}_{t}\in\Upsilon)\mathbb{P}_{K}^{-1}(\Upsilon)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{t}).$
Therefore, for any $\boldsymbol{\omega},\boldsymbol{\omega}^{\prime}\in\Omega$, by Lemma 19(ii) we can write
$KL\bigg(\tilde{\mathbf{P}}_{\boldsymbol{\omega},t},\tilde{\mathbf{P}}_{\boldsymbol{\omega}^{\prime},t}\bigg)=\int\log\left(\frac{\,\mathrm{d}\mathbf{P}_{\boldsymbol{\omega},t}}{\,\mathrm{d}\mathbf{P}_{\boldsymbol{\omega}^{\prime},t}}\right)\,\mathrm{d}\tilde{\mathbf{P}}_{\boldsymbol{\omega},t}$
$\phantom{00000000000}=\int\bigg[\log\left(\frac{\,\mathrm{d}F(y_{1}-f_{\boldsymbol{\omega}}(\boldsymbol{z}_{1},\boldsymbol{c}_{1}))}{\,\mathrm{d}F(y_{1}-f_{\boldsymbol{\omega}^{\prime}}(\boldsymbol{z}_{1},\boldsymbol{c}_{1}))}\right)+$
$\phantom{0000000}+\sum_{i=2}^{t}\log\left(\frac{\,\mathrm{d}F(y_{i}-f_{\boldsymbol{\omega}}\left(\pi_{i}\left(\boldsymbol{z}_{1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i}\right),\boldsymbol{c}_{i})\right)}{\,\mathrm{d}F(y_{i}-f_{\boldsymbol{\omega}^{\prime}}\left(\pi_{i}\left(\boldsymbol{z}_{1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i}\right),\boldsymbol{c}_{i})\right)}\right)\bigg]$
$\,\mathrm{d}F\left(y_{1}-f_{\boldsymbol{\omega}}\left(\boldsymbol{z}_{1},\boldsymbol{c}_{1}\right)\right)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{1})\left(\prod_{i=2}^{t-1}\,\mathrm{d}F\left(y_{i}-f_{\boldsymbol{\omega}}\left(\pi_{i}\left(\boldsymbol{z}_{1},y_{1}^{i-1},\boldsymbol{c}_{1}^{i}\right),\boldsymbol{c}_{i}\right)\right)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i})\right)$
$\phantom{00000000000000000}\,\mathrm{d}F\left(y_{t}-f_{\boldsymbol{\omega}}\left(\pi_{t}\left(\boldsymbol{z}_{1},y_{1}^{t-1},\boldsymbol{c}_{1}^{t}\right),\boldsymbol{c}_{t}\right)\right)\mathbb{1}(\boldsymbol{c}_{t}\in\Upsilon)\mathbb{P}_{K}^{-1}(\Upsilon)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{t})$
$\leq I_{0}\sum_{i=1}^{t-1}\int\max_{\boldsymbol{x}\in\Theta}|f_{\boldsymbol{\omega}}(\boldsymbol{x},\boldsymbol{c}_{i})-f_{\boldsymbol{\omega}^{\prime}}(\boldsymbol{x},\boldsymbol{c}_{i})|^{2}\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{i})$
$\phantom{00000}+I_{0}\int\max_{\boldsymbol{x}\in\Theta}|f_{\boldsymbol{\omega}}(\boldsymbol{x},\boldsymbol{c}_{t})-f_{\boldsymbol{\omega}^{\prime}}(\boldsymbol{x},\boldsymbol{c}_{t})|^{2}\mathbb{1}(\boldsymbol{c}_{t}\in\Upsilon)\mathbb{P}_{K}^{-1}(\Upsilon)\,\mathrm{d}\mathbb{P}_{K}(\boldsymbol{c}_{t}).$
Lemma 21
With the same notation as in the proof Theorem 11, let $f_{\boldsymbol{\omega},\boldsymbol{\tau}}:\mathbb{R}^{d}\times[0,1]^{p}\to\mathbb{R}$ denoted the function that it is defined in (42). Then, for all $\boldsymbol{\omega}\in\Omega$ and $\boldsymbol{\tau}\in\{-1,1\}^{d-1}$, $f_{\boldsymbol{\omega},\boldsymbol{\tau}}\in\mathcal{F}_{\alpha,\beta}(M)\cap\mathcal{F}_{\gamma}(L).$
Proof Fix $\boldsymbol{\omega}\in\Omega,\boldsymbol{\tau}\in\{-1,1\}^{d-1}$. First, let us prove that for all $\boldsymbol{x}\in\mathbb{R}^{d}$, $f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\cdot)$ is a $1$-Lipschitz function. For any $\boldsymbol{c},\boldsymbol{c}^{\prime}\in[0,1]^{p}$, we can write
$|f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})-f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}^{\prime})|$ $=r_{1}L\eta\left(x_{1}\delta^{-\frac{\gamma}{2}}\right)\bigg|\sum_{j=1}^{K^{p}}\omega_{j}\left(d^{\gamma}(\boldsymbol{c},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}\in B_{j}\right)-d^{\gamma}(\boldsymbol{c}^{\prime},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}^{\prime}\in B_{j}\right)\right)\bigg|$
$\leq r_{1}L\eta\left(1\right)\bigg|\sum_{j=1}^{K^{p}}\omega_{j}\left(d(\boldsymbol{c},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}\in B_{j}\right)-d^{\gamma}(\boldsymbol{c}^{\prime},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}^{\prime}\in B_{j}\right)\right)\bigg|.$
For $i,j\in K^{p}$, let us assume that $\boldsymbol{c}\in B_{i}$ and $\boldsymbol{c}^{\prime}\in B_{j}$, and let $\boldsymbol{y}_{i}\in\partial B_{i},\boldsymbol{y}^{\prime}_{j}\in\partial B_{j}$, such that $d(\boldsymbol{c},\partial B_{i})=\left\lVert\boldsymbol{c}-\boldsymbol{y}_{i}\right\rVert$ and $d(\boldsymbol{c}^{\prime},\partial B_{j})=\left\lVert\boldsymbol{c}^{\prime}-\boldsymbol{y}^{\prime}_{j}\right\rVert$. If $i=j$, then
$|f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})-f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}^{\prime})|$ $\leq r_{1}L\eta\left(1\right)\bigg|d^{\gamma}(\boldsymbol{c},\partial B_{i})-d^{\gamma}(\boldsymbol{c}^{\prime},\partial B_{i})\bigg|$
$=r_{1}L\eta\left(1\right)\bigg(\mathbb{1}\left(d(\boldsymbol{c},\partial B_{i})\geq d\left(\boldsymbol{c}^{\prime},\partial B_{i}\right)\right)\left(\left\lVert\boldsymbol{c}-\boldsymbol{y}_{i}\right\rVert-\left\lVert\boldsymbol{c}^{\prime}-\boldsymbol{y}^{\prime}_{i}\right\rVert\right)$
$\phantom{00000}+\mathbb{1}\left(d(\boldsymbol{c},\partial B_{i})<d\left(\boldsymbol{c}^{\prime},\partial B_{i}\right)\right)\left(\left\lVert\boldsymbol{c}^{\prime}-\boldsymbol{y}^{\prime}_{i}\right\rVert^{\gamma}-\left\lVert\boldsymbol{c}-\boldsymbol{y}_{i}\right\rVert^{\gamma}\right)\bigg)$
$\leq r_{1}L\eta\left(1\right)\bigg(\mathbb{1}\left(d(\boldsymbol{c},\partial B_{i})\geq d\left(\boldsymbol{c}^{\prime},\partial B_{i}\right)\right)\left(\left\lVert\boldsymbol{c}-\boldsymbol{y}^{\prime}_{i}\right\rVert^{\gamma}-\left\lVert\boldsymbol{c}^{\prime}-\boldsymbol{y}^{\prime}_{i}\right\rVert^{\gamma}\right)$
$\phantom{00000}+\mathbb{1}\left(d(\boldsymbol{c},\partial B_{i})<d\left(\boldsymbol{c}^{\prime},\partial B_{i}\right)\right)\left(\left\lVert\boldsymbol{c}^{\prime}-\boldsymbol{y}_{i}\right\rVert^{\gamma}-\left\lVert\boldsymbol{c}-\boldsymbol{y}_{i}\right\rVert^{\gamma}\right)\bigg)\leq L\left\lVert\boldsymbol{c}-\boldsymbol{c}^{\prime}\right\rVert^{\gamma},$
where the last inequality is obtained by the fact that $r_{1}\leq 1/(2\eta(1)\max(1,L))$. If $i\neq j$, it means that $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ does not belong to the same grid and we have
$d(\boldsymbol{c},\partial B_{i})\leq\left\lVert\boldsymbol{c}-\boldsymbol{c}^{\prime}\right\rVert,\quad\quad\text{and}\quad\quad d(\boldsymbol{c}^{\prime},\partial B_{j})\leq\left\lVert\boldsymbol{c}-\boldsymbol{c}^{\prime}\right\rVert.$
Consequently, we can write
$|f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})-f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c}^{\prime})|$ $\leq r_{1}L\eta\left(1\right)\left(d^{\gamma}\left(\boldsymbol{c},\partial B_{i}\right)+d^{\gamma}\left(\boldsymbol{c}^{\prime},\partial B_{i}\right)\right)\leq L\left\lVert\boldsymbol{c}-\boldsymbol{c}^{\prime}\right\rVert^{\gamma},$
where the last inequality is due to $r_{1}\leq 1/(2\eta(1)\max(1,L)).$
Fixing $\boldsymbol{c}\in[0,1]^{p}$, we calculate $\nabla_{\boldsymbol{x}}^{2}f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\cdot,\boldsymbol{c})$. Let $\left(\nabla^{2}_{\boldsymbol{x}}f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})\right)_{i,j}$ indicate the $(i,j)$ entry of the Hessian matrix. For $1\leq i\neq j\leq K^{p}$ it is straightforward to check that
$\left(\nabla^{2}_{\boldsymbol{x}}f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})\right)_{i,j}=0.$
Now, for $i=j=1$, we have
$|\left(\nabla^{2}_{\boldsymbol{x}}f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})\right)_{1,1}-2\alpha|\leq r_{1}L\delta^{-\gamma}\max_{x\in\mathbb{R}}|\nabla^{2}_{x}\eta\left(x\right)|\sum_{j=1}^{K^{p}}d^{\gamma}(\boldsymbol{c},\partial B_{j})\mathbb{1}\left(\boldsymbol{c}\in B_{j}\right)\leq\alpha,$
for $r_{1}\leq\alpha/\left(L^{\prime}\max\left(1,L\right)\right),$ where we introduced $L^{\prime}=\max_{x\in\mathbb{R}}|\eta^{\prime\prime}\left(x\right)|$. Furthermore, for $i=j>1$, we can write
$|\left(\nabla^{2}_{\boldsymbol{x}}f_{\boldsymbol{\omega},\boldsymbol{\tau}}(\boldsymbol{x},\boldsymbol{c})\right)_{i,i}-2\alpha|\leq r_{2}\max_{x\in\mathbb{R}}|\nabla^{2}_{x}\eta\left(x\right)|\leq\alpha,$
for $r_{2}\leq\alpha/L^{\prime}$. Finally, to show that $|f(\boldsymbol{x},\boldsymbol{c})|\leq M$ for all $\boldsymbol{x}\in\Theta$ and $\boldsymbol{c}\in[0,1]^{p}$, we have
$|f(\boldsymbol{x},\boldsymbol{c})|\leq\alpha+r_{1}L\eta(1)+r_{2}\eta(1)\leq M.$
Lemma 22
For any $\boldsymbol{c}\in B_{j}$ we have
$d(\boldsymbol{c},\partial B_{j})\leq\frac{1}{2K},$ (46)
and, for any $\varepsilon\in(0,1)$,
$\left\lVert\boldsymbol{c}-\boldsymbol{b}_{j}\right\rVert_{\infty}\leq\frac{\varepsilon}{2K}\quad\Longrightarrow\quad d(\boldsymbol{c},\partial B_{j})\geq\frac{1-\varepsilon}{2K},$ (47)
where $\left\lVert\cdot\right\rVert_{\infty}$ is the $\ell_{\infty}$ norm.
Proof Without loss of generality, we assume throughout the proof that $K=1$ and $\boldsymbol{b}_{j}=0$, so that the closure of $B_{j}$ is $1/2,1/2]^{d}$ and $\partial B_{j}=\{\boldsymbol{u}:\left\lVert\boldsymbol{u}\right\rVert_{\infty}=1/2\}$. Inequality (46) follows from the fact that $d(\boldsymbol{c},\partial B_{j})\leq d(\boldsymbol{c},\boldsymbol{c}^{\prime})\leq{1}/{2}$, where $\boldsymbol{c}^{\prime}\in\partial B_{j}$ is a vector such that all its coordinates except the first one coincide with those of $\boldsymbol{c}$, and the first coordinate $c^{\prime}_{1}$ of $\boldsymbol{c}^{\prime}$ is equal to the projection of the first coordinate $c_{1}\in[-1/2,1/2]$ of $\boldsymbol{c}$ onto the set $\{-1/2,1/2\}$.
In order to prove (47), we denote by $\boldsymbol{u}^{*}$ a solution of the problem
$\min_{\boldsymbol{u}:\left\lVert\boldsymbol{u}\right\rVert_{\infty}=1/2}\left\lVert\boldsymbol{c}-\boldsymbol{u}\right\rVert.$
There exists $i^{*}$ such that $|u_{i^{*}}^{*}|=1/2$, where $u_{i^{*}}^{*}$ is the $i^{*}$th coordinate of $\boldsymbol{u}^{*}$. Under the assumption that $\left\lVert\boldsymbol{c}\right\rVert_{\infty}\leq\varepsilon/2$ we have
$d^{2}(\boldsymbol{c},\partial B_{j})=\left\lVert\boldsymbol{c}-\boldsymbol{u}^{*}\right\rVert^{2}\geq|c_{i^{*}}-u_{i^{*}}^{*}|^{2}\geq\min(|c_{i^{*}}-1/2|^{2},|c_{i^{*}}+1/2|^{2})\geq\Big(\frac{1-\varepsilon}{2}\Big)^{2}.$
This proves (47).
Appendix D General comments on self-concordant barriers
In (Nesterov and Nemirovskii, 1994, Proposition 2.3.6), the authors demonstrated the existence of a convex body $\Theta$ such that any $\mu$-self-concordant barrier over $\Theta$ should have $\mu\geq d$. The proposition is stated as follows:
Statement of (Nesterov and Nemirovskii, 1994, Proposition 2.3.6): Let $\Theta\in\mathbb{R}^{d}$ be a convex polytope, such that certain boundary points of $\Theta$ belong to $m$ of (d1)(d-1)-dimensional facets of $\Theta$, where $m\in[d]$. Moreover, assume that the normal vectors to these facets are linearly independent. Then the value of $\mu$ for any $\mu$-self-concordant barrier on $\Theta$ cannot be less than $m$. In particular, the $d$-dimensional non-negative orthant, simplex, and hyper-cube do not admit barriers with the parameter less than $\mu$.
For the case when $\Theta$ is a polytope, one can construct the following self-concordant barrier (Nemirovski, 2004, Corollary 3.1.1)
  • $\mathcal{R}(\boldsymbol{u})=-\sum_{j=1}^{m}\log(\langle\boldsymbol{a}_{j},\boldsymbol{u}\rangle-b_{j})$ is an $m$-self-concordant barrier for the polytope
    $\Theta=\{\boldsymbol{u}\in\mathbb{R}^{d}:\langle\boldsymbol{a}_{j},\boldsymbol{u}\rangle\geq b_{j},\quad\text{for }j\in[m]\quad\text{and}\quad\boldsymbol{a}_{j}\in\mathbb{R}^{d},b_{j}\in\mathbb{R}\}.$
Note that in the above, if $\Theta$ is the $d$-dimensional hyper-cube then $m=2d$, and in this case (Nesterov and Nemirovskii, 1994, Proposition 2.3.6) indicates that the above $\mathcal{R}$ has an optimal $\mu$ as a function of $d$. Moreover, if $\Theta$ is the $d$-dimensional simplex then $m=d+1$, which again by (Nesterov and Nemirovskii, 1994, Proposition 2.3.6) we can see that the above example of $\mathcal{R}$ has an optimal $\mu$ as a function of $d$.
For the case when $\Theta$ is scaled version of the unit Euclidean ball with ratio $r>0$, one can consider the following self-concordant barrier (Nemirovski, 2004, Example 9.3.2)
  • $\mathcal{R}(\boldsymbol{u})=-\log(r^{2}-\left\lVert\boldsymbol{u}\right\rVert^{2})$ is a $2$-self-concordant barrier for $\Theta$.
Considering the result in (Nesterov and Nemirovskii, 1994, Corollary 2.3.3), we have always that $\mu\geq 1$ which shows the optimality of the above $\mathcal{R}$ for the unit Euclidean ball.
We refer the reader to (Nemirovski, 2004, Chapter 9) for further examples of self-concordant barrier functions.
In the general case of a convex body $\Theta$, Bubeck and Eldan (2014); Hildebrand (2014); Fox (2015), proposed $\mu$-self-concordant barriers with $\mu\leq\texttt{A}d$, where $\texttt{A}>0$ is a non-increasing function of $d$. Particularly, (Bubeck and Eldan, 2014, Theorem 1) shows that $\texttt{A}\leq 1+100\sqrt{\log(d)/d}$, for all $d\geq 80$. All of these constructions by Bubeck and Eldan (2014); Hildebrand (2014); Fox (2015) are computationally expensive (see (Bubeck and Eldan, 2014, Section 2.2) for further details on the computational aspects of these methods). Now, let us consider the special case where $\Theta$ is the unit $\ell_{q}$-ball for $q\in(2,\infty]$, i.e.,
$\Theta=\{\boldsymbol{u}\in\mathbb{R}^{d}:\sum_{j=1}^{d}|\boldsymbol{u}_{j}|^{q}\leq 1\},$
for $q\in(2,\infty)$, and
$\Theta=\{\boldsymbol{u}\in\mathbb{R}^{d}:\max_{j\in[d]}|\boldsymbol{u}_{j}|\leq 1\},$
for $q=\infty$. In (Bubeck et al., 2018, Lemma 11 and Theorem 5), the authors showed that any $\mu$-self-concordant barrier on $\Theta$ should satisfy
$\mu\geq\texttt{A}\max\left(d^{\frac{1}{q}}\log(d)^{-\frac{q}{q-2}},d^{1-\frac{2}{q}}\right),$
where $\texttt{A}>0$ only depends on $q$. The lower bound indicates that for the case of $q=\infty$, the $(d+1)$-self-concordant barrier proposed by Bubeck and Eldan (2014); Hildebrand (2014); Fox (2015) is optimal with respect to $\mu$ as a function of $d$. However, constructing an optimal $\mu$-self-concordant barrier for any $\ell_{q}$-ball for $q\in(2,\infty)$ remains an open problem.